### v2

# Ordinal graphs and their $\mathrm{C}^{*}$-algebras

Abstract.
We introduce a class of left cancellative categories we call ordinal graphs for which there is a functor $d:\Lambda\rightarrow\mathrm{Ord}$ by which morphisms of $\Lambda$ factor. We use generators and relations to study the Cuntz-Krieger algebra $\mathcal{O}\left(\Lambda\right)$ defined by Spielberg. In particular, we construct a $\mathrm{C}^{*}$-correspondence $X_{\alpha}$ for each $\alpha\in\mathrm{Ord}$ in order to apply Eryüzlü and Tomforde’s condition (S) and prove a Cuntz-Krieger uniqueness theorem for ordinal graphs.
Key words and phrases:
Ordinal, Graph, Cuntz-Pimsner algebra, $\mathrm{C}^{*}$-correspondence
2000 Mathematics Subject Classification:
Primary 46L05
1. Introduction
A directed graph $E$ is a collection $\left(E^{0},E^{1},r,s\right)$ where $E^{0}$, $E^{1}$ are countable, discrete sets and $r,s:E^{1}\rightarrow E^{0}$. To each directed graph $E$ there is an associated $\mathrm{C}^{*}$-algebra $C^{*}\left(E\right)$. This algebra is defined using generators and relations, as in [9, Chapter 5]. Namely, $C^{*}\left(E\right)$ is the $\mathrm{C}^{*}$-algebra which is universal for mutually orthogonal projections $\left\{P_{v}:v\in E^{0}\right\}$ and partial isometries $\left\{S_{e}:e\in E^{1}\right\}$ with mutually orthogonal ranges such that for $v\in E^{0}$ and $e\in E^{1}$,
  1. (1)
    $S_{e}^{*}S_{e}=P_{s(e)}$
  2. (2)
    $P_{r(e)}S_{e}S_{e}^{*}=S_{e}S_{e}^{*}$
  3. (3)
    $P_{v}=\sum_{f\in r^{-1}\left(v\right)}S_{f}S_{f}^{*}$ if $0<\left|r^{-1}\left(v\right)\right|<\infty$
Regarding $\mathbb{N}$ as a monoid under addition, one may equivalently define a graph as a small category $\Lambda$ with a functor $d:\Lambda\rightarrow\mathbb{N}$ with the following factorization property: for every $e\in\Lambda$ and $n\leq d\left(e\right)$ there are unique $f,g\in\Lambda$ with $d\left(f\right)=n$ and $e=fg$. This enables many generalizations of the construction of the algebra $C^{*}\left(E\right)$ above. In [8], Kumjian and Pask replace $\mathbb{N}$ with the monoid $\mathbb{N}^{k}$ for some $k\in\mathbb{N}$ to define $k$-graphs. Then a similar set of generators and relations yields an algebra $C^{*}\left(\Lambda\right)$ which recovers $C^{*}\left(E\right)$ in the case $k=1$.
More recently, Brown and Yetter in [2] replace $\mathbb{N}$ with a category $\mathcal{B}$ and use groupoids to derive conditions for Cuntz-Krieger uniqueness. In particular, they obtain Cuntz-Krieger uniqueness when the factorization satisfies certain conditions and $\mathcal{B}$ is right-cancellative. Lydia de Wolf further develops this theory and gives examples of this construction when $\mathcal{B}=\mathbb{N}_{0}^{\mathbb{N}}$ and $\mathcal{B}=\mathbb{Q}_{\geq 0}$ in [3]. One may also consider $C^{*}\left(\Lambda\right)$ for a category $\Lambda$ when $\Lambda$ has no such functor $d$ [11]. All that is required in this construction is a small category $\Lambda$ which is left-cancellative, yet in the case that $\Lambda$ is a directed graph or $k$-graph, the usual $\mathrm{C}^{*}$-algebra is recovered.
In this paper, we consider the case in which there is a functor $d:\Lambda\rightarrow\mathrm{Ord}$ with the factorization property, where $\mathrm{Ord}$ denotes the ordinals. We regard the ordinals as a monoid under ordinal addition and call the category $\Lambda$ with this functor an ordinal graph. The natural numbers are a submonoid of $\mathrm{Ord}$, which makes ordinal graphs a generalization of directed graphs. We use generators and relations from [11] to analyze $C^{*}\left(\Lambda\right)=\mathcal{O}\left(\Lambda\right)$. While addition on $\mathbb{N}$ is abelian and cancellative, addition on $\mathrm{Ord}$ is neither abelian nor right-cancellative, and this has consequences for the types of relations we may represent with ordinal graphs. Instead of groupoids, we adopt an approach using $\mathrm{C}^{*}$-correpsondences and apply results from [4] to recover a Cuntz-Krieger uniqueness result.
2. Preliminaries
If $\left(X,\phi\right)$ is a $\mathrm{C}^{*}$-correspondence over $A$, we write $a\cdot x$ for $\phi\left(a\right)x$. For each such correspondence there is a closed ideal $J_{X}\vartriangleleft A$ with $J_{X}=\left(\ker\phi\right)^{\perp}\cap\phi^{-1}\left(\mathcal{K}\left(X\right)\right)$ and an associated Cuntz-Pimsner algebra $\mathcal{O}\left(X\right)$ obtained as the universal $\mathrm{C}^{*}$-algebra for covariant representations [6]. Maps $\psi:X\rightarrow\mathcal{C}$ and $\pi:A\rightarrow\mathcal{C}$ into a $\mathrm{C}^{*}$-algebra $\mathcal{C}$ form a representation $\left(\psi,\pi\right)$ when $\psi$ is $A$-linear, $\pi$ is a *-homomorphism, $\psi\left(x\cdot a\right)=\psi\left(x\right)\pi\left(a\right)$, $\psi\left(x\right)^{*}\psi\left(y\right)=\pi\left(\left\langle x,y\right\rangle\right)$, and $\psi\left(a\cdot x\right)=\pi\left(a\right)\psi\left(x\right)$. Associated to each representation $\left(\psi,\pi\right)$ is a *-homomorphism $\left(\psi,\pi\right)^{\left(1\right)}:\mathcal{K}\left(X\right)\rightarrow\mathcal{C}$ where $\left(\psi,\pi\right)^{\left(1\right)}\left(\theta_{x,y}\right)=\psi\left(x\right)\psi\left(y\right)^{*}$. Then the representation $\left(\psi,\pi\right)$ is covariant if for every $a\in J_{X}$, $\left(\psi,\pi\right)^{\left(1\right)}\left(\phi\left(a\right)\right)=\pi\left(a\right)$. For each correspondence $\left(X,\phi\right)$ over $A$, one can construct correspondences $\left(X^{\otimes n},\phi_{n}\right)$ over $A$, where $X^{\otimes n}$ is the $n$-fold internal tensor product of $X$ over $A$. Given a representation $\left(\psi,\pi\right)$ of $\left(X,\phi\right)$, there exists representations $\left(\psi^{\otimes n},\pi\right)$ of $\left(X^{\otimes n},\phi_{n}\right)$ [5, Proposition 1.8].
Throughout the paper we will make use of ordinal arithmetic. If $\leq$ is a partial order on a set $A$, we say $\leq$ well-orders $A$ if for all $E\subseteq A$ with $E\not=\emptyset$, there exists $x\in E$ with $x=\min E$. An order isomorphism between two partially ordered sets is an increasing bijection $f:A\rightarrow B$ with an increasing inverse. An ordinal is an order isomorphism class of well-ordered sets. We denote the order isomorphism class of a well-ordered set $A$ as $\left[A\right]$.
Ordinal addition is defined so that $\left[A\right]+\left[B\right]=\left[A\sqcup B\right]$ where $A\sqcup B$ is well-ordered by a relation $\leq$ which agrees with the order on $A$ and $B$ such that $a\leq b$ for all $a\in A$ and $b\in B$. Addition induces a well order on Ord, where for $\alpha,\beta\in\text{Ord}$, $\alpha\leq\beta$ if and only if there is $\gamma\in\text{Ord}$ such that $\alpha+\gamma=\beta$. Similarly, ordinal multiplication is defined by $\left[A\right]\cdot\left[B\right]=\left[A\times B\right]$ where $\left(a,b\right)\leq\left(c,d\right)$ if $b<d$, or if $b=d$ and $a\leq c$. For ordinals $\left[A\right]$ and $\left[B\right]$, exponentiation $\left[A\right]^{\left[B\right]}$ is defined to be $\left[C\right]$, where
$C=\left\{f\in A^{B}:B\backslash f^{-1}\left(\min A\right)\text{ is finite}\right\}$
Here $A^{B}$ denotes the set of functions from $B$ to $A$. Then $C$ is well-ordered, so that for distinct $f,g\in C$ and $x=\max\left\{b\in B:f\left(b\right)\not=g\left(b\right)\right\}$, $f<g$ if and only if $f\left(x\right)<g\left(x\right)$ [10, Chapter XIV].
If $A$ is a finite well-ordered set, we say $\left[A\right]$ is finite. Every finite well-ordered set is order isomorphic to a bounded interval in $\mathbb{N}$, and for $n\in\mathbb{N}$, we identify the ordinal [$\left[\left\{k\in\mathbb{N}:k<n\right\}\right]$ with $n$. Furthermore, arithmetic for finite ordinals agrees with arithmetic on $\mathbb{N}$. We define $\omega=\left[\mathbb{N}\right]$, from which we see $\mathbb{N}$ is order isomorphic to $\left\{\alpha\in\text{Ord}:\alpha<\omega\right\}=\left[0,\omega\right)$.
While ordinal arithmetic and arithmetic on $\mathbb{N}$ share many useful properties, for infinite ordinals $\alpha,\beta\in\mathrm{Ord}$, it is no longer necessary that $\alpha+\beta=\beta+\alpha$ or $\alpha\cdot\beta=\beta\cdot\alpha$. However, we still have $\alpha\cdot\left(\beta+\gamma\right)=\alpha\cdot\beta+\alpha\cdot\gamma$. We also have a division algorithm for $\mathrm{Ord}$: if $\gamma<\alpha\cdot\beta$, there exist unique $\beta_{1}<\beta$ and $\alpha_{1}<\alpha$ such that $\gamma=\alpha\cdot\beta_{1}+\alpha_{1}$. Base expansion in terms of infinite bases makes sense for ordinals by the following theorem.
Theorem 2.1 ([10, Chapter XIV 19.3]).
For every $\alpha>0$ and $\beta>1$ there exists unique $k\in\mathbb{N}$ such that $\alpha$ may be represented uniquely as
$\alpha=\beta^{\alpha_{1}}\cdot\gamma_{1}+\beta^{\alpha_{2}}\cdot\gamma_{2}+\ldots+\beta^{\alpha_{k}}\cdot\gamma_{k}$
where $\alpha_{n}\geq\alpha_{n+1}$ and $0<\gamma_{n}<\beta$.
When $\beta=\omega$, 2.1 gives a representation of $\alpha$ named the Cantor normal form. Since in that case the coefficients $\gamma_{n}$ are finite, every ordinal $\alpha$ may be written as a finite sum of powers of $\omega$.
For ordinals $\alpha,\beta$ with $\alpha<\beta$, $\omega^{\alpha}+\omega^{\beta}=\omega^{\beta}$. Applying this fact and the Cantor normal form, it’s possible to compute arbitrary sums and products of ordinals using the algebraic properties of ordinals we have stated here. Moreover, addition of ordinals is left cancellative: for $\alpha,\beta,\gamma\in\mathrm{Ord}$, $\alpha+\beta=\alpha+\gamma$ implies $\beta=\gamma$. Thus we have one-sided subtraction of ordinals. If $\alpha\leq\beta$, we write $-\alpha+\beta$ for the unique ordinal $\gamma$ such that $\alpha+\gamma=\beta$. Additionally, we write the expression $\delta-\alpha+\beta$ for the ordinal $\delta+\left(-\alpha+\beta\right)$. Using this notation, we have $-\omega^{\alpha}+\omega^{\beta}=\omega^{\beta}$ when $\alpha<\beta$, and for $\alpha=0,\beta=1$, this gives $1+\omega=-1+\omega=\omega$.
3. Ordinal Graphs
We begin by defining ordinal graphs. The definition is almost identical to [8, Definition 1.1], except the range of $d$ is $\mathrm{Ord}$ instead of $\mathbb{N}^{k}$.
Definition 3.1.
An ordinal graph is a pair $\left(\Lambda,d\right)$ where $\Lambda$ is a small category and $d:\Lambda\rightarrow\mathrm{Ord}$ is a functor such that for every $e\in\Lambda$ and $\alpha\leq d(e)$, there exist unique $f,g\in\Lambda$ with $d(f)=\alpha$ and $e=fg$.
Here small category means $\Lambda$ is a set of morphisms. We will often suppress the functor $d$ and simply write $\Lambda$ for an ordinal graph $\left(\Lambda,d\right)$. We call $d$ the length functor, and elements of $\Lambda$ paths. $\mathrm{Ord}$ is regarded as a monoid under ordinal addition. Then since $d$ is a functor, $d\left(ef\right)=d\left(e\right)+d\left(f\right)$ for paths $e,f\in\Lambda$. We frequently refer to the existence and uniqueness of $f$ and $g$ in 3.1 as unique factorization.
It may happen that $d$ maps into the ordinals less than $\omega$. Then since the range of $d$ is order isomorphic to a subset of $\mathbb{N}$, the factorization property ensures that each path is a product of paths of length $1$. Hence $\Lambda$ is the category of paths of a directed graph $E$, where $E^{0}=d^{-1}\left(0\right)$ and $E^{1}=d^{-1}\left(1\right)$. Since infinite ordinals are not finite sums of the ordinal $1$, we do not in general have that $\Lambda$ is generated by $d^{-1}\left(1\right)$.
Definition 3.2.
Let $\left(\Lambda,d\right)$ be an ordinal graph. We call $v\in\Lambda$ a vertex if $d\left(v\right)=0$.
As a consequence of uniqueness in the factorization property, we obtain left cancellation.
Proposition 3.3.
Every ordinal graph is a left cancellative small category.
Proof.
Let $\left(\Lambda,d\right)$ be an ordinal graph, and suppose $e=fg=fh$ for $f,g,h\in\Lambda$. Then $fg$ and $fh$ are factorizations of $e$, so by unique factorization $g=h$. ∎
Remark 3.4.
It may be tempting to assume that the same argument yields right cancellation. This is not true, however, because the factorization of $e$ into $fg$ in 3.1 only allows us to control the length of $f$. More precisely, $e$ may factor as $fg=hk$ where $d\left(g\right)=d\left(k\right)$ and $f\not=h$. The following example illustrates this.
Example 3.5.
Let $\Lambda=\left[0,\omega^{2}\right)$ and $d:\Lambda\rightarrow\mathrm{Ord}$ be the restriction of the identity function. Then $\Lambda$ is a category with composition given by ordinal addition. This follows because for every $\alpha,\beta\in\Lambda$, $\alpha<\omega^{2}$ and $\beta<\omega^{2}$, so $\alpha+\beta<\omega^{2}$ and $\Lambda$ is closed under composition. Unique factorization follows from ordinal subtraction. In particular, for every $\alpha\in\Lambda$ and $\beta\leq d\left(\alpha\right)=\alpha$ there exists unique $\gamma\leq d\left(\alpha\right)$ such that $\beta+\gamma=\alpha$. Thus $\left(\Lambda,d\right)$ forms an ordinal graph.
$v$$e$$f:eee\ldots;ef=f$
Figure 3.1. $\Lambda=\left[0,\omega^{2}\right)$ as an ordinal graph, where $v=0$, $e=1$, and $f=\omega$
To distinguish between paths in $\Lambda$ and elements of $\mathrm{Ord}$, we may define $v=0\in\Lambda$, $e=1\in\Lambda$, and $f=\omega\in\Lambda$. Recall that each ordinal can be written uniquely as a sum of powers of $\omega$. Then $vf=ef=f$, and thus every path in $\Lambda$ is of the form $\underbrace{f\ldots f}_{n\text{ times}}\underbrace{e\ldots e}_{m\text{ times}}v$.
In Fig. 3.1, we write $f:eee\ldots$ to denote a path $f$ which factors as $f=eeeg$ for some path $g$. In general, we write $h:e_{1}e_{2}e_{3}\ldots$ for a path $h$ which for each $k\in\mathbb{N}$ factors as $h=e_{1}e_{2}\ldots e_{k}g_{k}$ for some paths $g_{k}$ and satisfies
$d\left(h\right)=\sup_{n\in\mathbb{N}}\sum_{k=1}^{n}d\left(e_{k}\right)$
Indeed, in this example we have $e_{k}=e$ and $h=f$. Since $ef=f$, the notation is redundant, but it is useful in the next example.
Example 3.6.
Define a category $\Lambda$ generated by $\left\{v_{n},e_{n},f_{n},g_{n}:n\in\mathbb{N}\right\}$ with relations:
  1. (1)
    $v_{n}=r\left(e_{n}\right)=r\left(f_{n}\right)=s\left(f_{n}\right)=r\left(g_{n}\right)$
  2. (2)
    $v_{n}=s\left(e_{n-1}\right)$ for $n>0$
  3. (3)
    $g_{n}=f_{n}e_{n}g_{n+1}$
  4. (4)
    $s\left(g_{0}\right)=v_{0}$
Then $\left\{v_{n}:n\in\mathbb{N}\right\}$ are the identity maps of $\Lambda$. There is a functor $d:\Lambda\rightarrow\mathrm{Ord}$ with $d\left(v_{n}\right)=0$, $d\left(e_{n}\right)=d\left(f_{n}\right)=1$, and $d\left(g_{n}\right)=\omega$. We may regard elements of $\Lambda$ as finite compositions of the generators. Then if $e,f,g\in\Lambda$ are distinct with $d\left(g\right)=\omega$, $d\left(e\right)=d\left(f\right)=1$, and $s\left(e\right)=s\left(f\right)=r\left(g\right)$, we have $eg=fg$ only if $e=f$. This ensures $\Lambda$ has the unique factorization property.
$v_{0}$$v_{1}$$v_{2}$$f_{0}$$f_{1}$$f_{2}$$e_{2}$$e_{1}$$e_{0}$$g_{2}:f_{2}e_{2}f_{3}e_{3}\dots$
Figure 3.2. The ordinal graph $\Lambda$ in Example 3.6
By 3.3 and [11, 10.3], each ordinal graph has an associated Toeplitz $\mathrm{C}^{*}$-algebra $\mathcal{T}(\Lambda)$ and a Cuntz-Krieger $\mathrm{C}^{*}$-algebra $\mathcal{O}(\Lambda)$. We will work with these algebras using generators and relations, which we record later in this section.
Definition 3.7.
If $\Lambda$ is an ordinal graph and $\alpha\in\mathrm{Ord}$, define the ordinal graph $\left(\Lambda_{\alpha},d_{\alpha}\right)$ by $\Lambda_{\alpha}=\left\{e\in\Lambda:d(e)<\omega^{\alpha}\right\}$ and $d_{\alpha}=\left.d\right|_{\Lambda_{\alpha}}$. Define $\Lambda^{\alpha}=\left\{e\in\Lambda:d(e)=\omega^{\alpha}\right\}$. The elements $v$ of $\Lambda_{0}$ are the vertices of $\Lambda$.
Note that $\Lambda_{0}=\Lambda^{0}$. $\Lambda_{\alpha}$ is well-defined because whenever $\beta,\gamma<\omega^{\alpha}$, $\beta+\gamma<\omega^{\alpha}$. From $\Lambda$, we may construct $\mathcal{T}\left(\Lambda_{\alpha}\right)$ and $\mathcal{O}\left(\Lambda_{\alpha}\right)$ for all ordinals $\alpha$. Since $\Lambda$ is a set and $\mathrm{Ord}$ is a proper class, $\mathcal{T}\left(\Lambda_{\alpha}\right)=\mathcal{T}\left(\Lambda\right)$ and $\mathcal{O}\left(\Lambda_{\alpha}\right)=\mathcal{O}\left(\Lambda\right)$ for sufficiently large $\alpha\in\mathrm{Ord}$. Later we will see there is much to say about the relationship between these algebras.
Since $\Lambda$ is a category, it comes equipped with source and range maps $s,r:\Lambda\rightarrow\Lambda$. As with directed graphs, $s\left(e\right)$ and $r\left(e\right)$ should be interpreted as the source and range vertices of the path $e$.
Proposition 3.8.
$v\in\Lambda$ is a vertex if and only if $v=s\left(e\right)$ for some $e\in\Lambda$.
Proof.
Suppose $v=s\left(e\right)$. Then $vv=s\left(e\right)s\left(e\right)=s\left(e\right)=v$, hence $d\left(v\right)+d\left(v\right)=d\left(v\right)$. Subtracting $d\left(v\right)$, we see $d\left(v\right)=0$. Since $s\left(r\left(e\right)\right)=r\left(e\right)$, this argument also implies $r\left(e\right)\in\Lambda_{0}$. On the other hand, if $d\left(v\right)=0$, then $r\left(v\right)vs\left(v\right)=v$. Since $d\left(r\left(v\right)v\right)=d\left(r\left(v\right)\right)=0$, unique factorization implies $v=r\left(v\right)v=r\left(v\right)$, and $s\left(v\right)=s\left(r\left(v\right)\right)=r\left(v\right)=v$. ∎
Definition 3.9.
For $e\in\Lambda$ with $\alpha\leq d(e)$, write $e_{\alpha}$ and $e^{\alpha}$ for the unique paths satisfying $e_{\alpha}e^{\alpha}=e$ and $d\left(e_{\alpha}\right)=\alpha$.
For finite ordinals $\alpha$, $e^{\alpha}$ may be confused with a power of $e$. We will never use exponents to denote repeated composition. We must also take care to avoid confusing $e_{\alpha}$ with a member of a sequence of paths $e_{1},e_{2},e_{3},\ldots$ For finite $\alpha$ and a path $e\in\Lambda$, $e_{\alpha}$ will always refer to the factorization of $e$. If we wish to denote a sequence of paths $f_{1},f_{2},\ldots$, then we will avoid ambiguity by not labeling any path in $\Lambda$ as $f$.
We record some calculations that apply to the notation in the previous definition. We frequently make use of these identities.
Proposition 3.10.
Let $e,f\in\Lambda$ with $s\left(e\right)=r\left(f\right)$. Then
  1. a.
    If $\alpha\leq d\left(e\right)$, $\left(ef\right)_{\alpha}=e_{\alpha}$ and $\left(ef\right)^{\alpha}=e^{\alpha}f$.
  2. b.
    If $\alpha\geq d\left(e\right)$, then $\left(ef\right)_{\alpha}=ef_{-d\left(e\right)+\alpha}$ and $\left(ef\right)^{\alpha}=f^{-d\left(e\right)+\alpha}$.
  3. c.
    If $\alpha+\beta\leq d\left(e\right)$, then $\beta\leq d\left(e^{\alpha}\right)$. Moreover, $\left(e^{\alpha}\right)^{\beta}=e^{\alpha+\beta}$ and $\left(e^{\alpha}\right)_{\beta}=\left(e_{\alpha+\beta}\right)^{\alpha}$.
  4. d.
    If $\beta\leq\alpha\leq d\left(e\right)$, $\left(e_{\alpha}\right)^{\beta}=\left(e^{\beta}\right)_{-\beta+\alpha}$ and $\left(e_{\alpha}\right)_{\beta}=e_{\beta}$.
Proof.
For part a, suppose $\alpha\leq d\left(e\right)$. Note that $d\left(e_{\alpha}\right)=\alpha$ and $e_{\alpha}e^{\alpha}f=ef$, so by unique factorization, $\left(ef\right)_{\alpha}=e_{\alpha}$ and $\left(ef\right)^{\alpha}=e^{\alpha}f$. For b, suppose on the other hand that $\alpha\geq d\left(e\right)$. Then $d\left(ef_{-d\left(e\right)+\alpha}\right)=d\left(e\right)+d\left(f_{-d\left(e\right)+\alpha}\right)=d\left(e\right)-d\left(e\right)+\alpha=\alpha$, and $ef=ef_{-d\left(e\right)+\alpha}f^{-d\left(e\right)+\alpha}$. Unique factorization once again gives the desired equalities. For part c, assume $\alpha+\beta\leq d\left(e\right)$. Then $\alpha+\beta\leq d\left(e\right)=d\left(e_{\alpha}e^{\alpha}\right)=d\left(e_{\alpha}\right)+d\left(e^{\alpha}\right)=\alpha+d\left(e^{\alpha}\right)$. Subtracting $\alpha$ from both sides, we see $\beta\leq d\left(e^{\alpha}\right)$. Also, $e_{\alpha}\left(e^{\alpha}\right)_{\beta}\left(e^{\alpha}\right)^{\beta}=e=e_{\alpha+\beta}e^{\alpha+\beta}$, and $d\left(e_{\alpha}\left(e^{\alpha}\right)_{\beta}\right)=\alpha+\beta=d\left(e_{\alpha+\beta}\right)$. By unique factorization, $e_{\alpha}\left(e^{\alpha}\right)_{\beta}=e_{\alpha+\beta}$ and $\left(e^{\alpha}\right)^{\beta}=e^{\alpha+\beta}$. Then since $d\left(e_{\alpha}\right)=\alpha$, $\left(e_{\alpha+\beta}\right)^{\alpha}=\left(e^{\alpha}\right)_{\beta}$. Finally, assume for part d that $\alpha\leq\beta\leq d\left(e\right)$. $\left(e_{\alpha}\right)^{\beta}=\left(e^{\beta}\right)_{-\beta+\alpha}$ follows from part c, and moreover $\left(e_{\alpha}\right)_{\beta}\left(e_{\alpha}\right)^{\beta}e^{\alpha}=e=e_{\beta}e^{\beta}$ with $d\left(\left(e_{\alpha}\right)_{\beta}\right)=\beta=d\left(e_{\beta}\right)$. Hence $\left(e_{\alpha}\right)_{\beta}=e_{\beta}$ by unique factorization.
If $\Lambda$ is a left cancellative small category which is finitely aligned in the following sense, then $\mathcal{T}\left(\Lambda\right)$ and $\mathcal{O}\left(\Lambda\right)$ may be understood in terms of generators, relations, and the exhaustive sets of $\Lambda$. We record the relevant definitions below. ∎
Definition 3.11 ([11, Definition 3.2]).
A left cancellative category $\Lambda$ is finitely aligned if for all $e,f\in\Lambda$ there exists finite $F\subseteq\Lambda$ such that $e\Lambda\cap f\Lambda=\cup_{g\in F}g\Lambda$.
Definition 3.12 ([11, Definition 10.13]).
If $\Lambda$ is a left cancellative category and $v\in\Lambda$, $F\subseteq v\Lambda$ is exhaustive for $v$ if for all $f\in v\Lambda$ there exists $e\in F$ such that $e\Lambda\cap f\Lambda\not=\emptyset$.
Lemma 3.13.
Every ordinal graph $\Lambda$ is finitely aligned. Moreover, if $e,f\in\Lambda$, then $e\Lambda\cap f\Lambda\in\left\{\emptyset,e\Lambda,f\Lambda\right\}$.
Proof.
Let $e,f\in\Lambda$ be given, and suppose without loss of generality that $d(e)\leq d(f)$. Choose $p,q\in\Lambda$ such that $f=pq$ with $d(p)=d(e)$. If $e\Lambda\cap f\Lambda=\emptyset$, then we are done. Otherwise, let $g\in e\Lambda\cap f\Lambda$ and choose $r,s\in\Lambda$ such that $g=er=ps$. By unique factorization, $p=e$, and therefore $e\Lambda\cap f\Lambda=e\Lambda\cap eq\Lambda=e\Lambda$. ∎
To present $\mathcal{T}\left(\Lambda\right)$ and $\mathcal{O}\left(\Lambda\right)$ in terms of generators and relations, we must study the common extensions in $\Lambda$.
Definition 3.14 ([11, Definition 2.4]).
For a left cancellative category $\Lambda$ and $e,f\in\Lambda$, $e\approx f$ if there is invertible $g\in\Lambda$ with $eg=f$.
Definition 3.15 ([11, Definition 3.1]).
If $F$ is a subset of a left cancellative category $\Lambda$, the common extensions of $F$ are the elements of $\bigcap_{e\in F}e\Lambda$. If $e$ is a common extension of $F$, $e$ is minimal if for all common extensions $f$ of $F$ with $e\in f\Lambda$, $e\approx f$. $\bigvee F$ denotes the set of minimal common extensions of $F$. For $e,f\in\Lambda$, $e\vee f$ denotes $\bigvee\{e,f\}$.
Lemma 3.16.
If $\Lambda$ is an ordinal graph and $e,f\in\Lambda$, $e\approx f$ if and only if $e=f$.
Proof.
First suppose $e\approx f$. Choose invertible $g\in\Lambda$ with $eg=f$. Then $d\left(s\left(g\right)\right)=d\left(s\left(g\right)s\left(g\right)\right)=d\left(s\left(g\right)\right)+d\left(s\left(g\right)\right)$, hence $d\left(s\left(g\right)\right)=0$. Also, $d\left(g^{-1}g\right)=d\left(s\left(g\right)\right)=d\left(g^{-1}\right)+d\left(g\right)=0$, so $d\left(g\right)=0$. Thus $d\left(eg\right)=d\left(e\right)+d\left(g\right)=d\left(e\right)=d\left(f\right)$, and by unique factorization of $eg=fs\left(f\right)$, $e=f$. Clearly if $e=f$, $es\left(e\right)=f$ and $e\approx f$. ∎
Lemma 3.17.
If $\Lambda$ is an ordinal graph, $e,f\in\Lambda$, and $e\Lambda\cap f\Lambda\not=\emptyset$, then either $e\in f\Lambda$ or $f\in e\Lambda$. If $e\in f\Lambda$, $e\vee f=\left\{f\right\}$, and if $f\in e\Lambda$, $e\vee f=\left\{e\right\}$.
Proof.
Let $e,f\in\Lambda$ with $e\Lambda\cap f\Lambda\not=\emptyset$. By 3.13, either $e\Lambda\cap f\Lambda=e\Lambda$ or $e\Lambda\cap f\Lambda=f\Lambda$. We may assume without loss of generality that $e\Lambda\cap f\Lambda=e\Lambda$. Hence the common extensions of $\left\{e,f\right\}$ are the elements of $e\Lambda$. If $g$ is a common extension of $\left\{e,f\right\}$ and $e\in g\Lambda$, then $g=eh$ for some $h\in\Lambda$, and $e\in eh\Lambda$. Thus $d\left(e\right)\geq d\left(e\right)+d\left(h\right)$, and $d\left(h\right)=0$. Therefore $g=e$, and $e$ is minimal. If $g\in e\Lambda$ is minimal, then since $e$ is a common extension, $e\approx g$, and by 3.16, $g=e$. We conclude that $e\vee f=\left\{e\right\}$. ∎
We apply the following theorems to obtain generators and relations for $\mathcal{T}\left(\Lambda\right)$ and $\mathcal{O}\left(\Lambda\right)$.
Theorem 3.18 ([11, Theorem 10.15], [11, Theorem 9.7]).
Let $\Lambda$ be a countable finitely-aligned left cancellative small category. Then $\mathcal{O}\left(\Lambda\right)$ is the $C^{*}$-algebra which is universal for generators $\left\{T_{e}:e\in\Lambda\right\}$ and relations 1-4. $\mathcal{T}\left(\Lambda\right)$ is the $C^{*}$-algebra which is universal for generators $\left\{T_{e}:e\in\Lambda\right\}$ and relations 1-3.
  1. (1)
    $T_{e}^{*}T_{e}=T_{s(e)}$
  2. (2)
    $T_{e}T_{f}=T_{ef}$ if $s(e)=r(f)$
  3. (3)
    $T_{e}T_{e}^{*}T_{f}T_{f}^{*}=\bigvee_{g\in e\vee f}T_{g}T_{g}^{*}$
  4. (4)
    $T_{v}=\bigvee_{e\in F}T_{e}T_{e}^{*}$ if $F$ is finite and exhaustive for $v$
Note that by 3.17, relation 3 for ordinal graphs may be rewritten as
  1. (3)
    $T_{e}T_{e}^{*}T_{f}T_{f}^{*}=\begin{cases}T_{e}T_{e}^{*}&e\in f\Lambda\\ T_{f}T_{f}^{*}&f\in e\Lambda\\ 0&\text{otherwise}\end{cases}$
Even when an ordinal graph $\Lambda$ is not countable, we define $\mathcal{O}\left(\Lambda\right)$ and $\mathcal{T}\left(\Lambda\right)$ using these relations. If $\mathcal{C}$ is a $\mathrm{C}^{*}$-algebra and $T=\left\{T_{e}:e\in\Lambda\right\}\subseteq\mathcal{C}$ is a family of operators satisfying relations 1-4, then we call $T$ a Cuntz-Krieger $\Lambda$-family. If instead $T$ satisfies relations 1-3, we call $T$ a Toeplitz $\Lambda$-family. Every Toeplitz $\Lambda$-family $T\subseteq\mathcal{C}$ induces a *-homomorphism $\pi:\mathcal{T}\left(\Lambda\right)\rightarrow\mathcal{C}$ mapping generators to members of $T$. Likewise, each Cuntz-Krieger family induces a *-homomorphism $\pi:\mathcal{O}\left(\Lambda\right)\rightarrow\mathcal{C}$.
We record a few calculations which will prove useful in the study of these $\mathrm{C}^{*}$-algebras.
Lemma 3.19.
Let $\Lambda$ be an ordinal graph and $e,f\in\Lambda$. The following hold in $\mathcal{T}\left(\Lambda\right)$ and $\mathcal{O}\left(\Lambda\right)$.
  1. a.
    $T_{e}$ is a partial isometry.
  2. b.
    If $v\in\Lambda_{0}$, then $T_{v}$ is a projection.
  3. c.
    If $s(e)\not=r(f)$, then $T_{e}T_{f}=0$.
  4. d.
    $T_{e}^{*}T_{f}=\begin{cases}T_{g}&f=eg\\ T_{g}^{*}&e=fg\\ 0&e\Lambda\cap f\Lambda=\emptyset\end{cases}$
Proof.
For part a, notice that $s(e)=s(s(e))$, so by relation 2, $T_{e}T_{s(e)}=T_{es(e)}=T_{e}$. Then by relation 1, $T_{e}T_{e}^{*}T_{e}=T_{e}T_{s(e)}=T_{e}$, and thus $T_{e}$ is a partial isometry. For part b, we have $s(v)=v$, and relation 1 implies $T_{v}^{*}T_{v}=T_{v}$. To prove part c, suppose $s(e)\not=r(f)$. Then $s(e)\Lambda\cap r(f)\Lambda=\emptyset$, and by relation 3, $T_{s(e)}T_{s(e)}^{*}T_{r(f)}T_{r(f)}^{*}=0$. Since $T_{s(e)}$ and $T_{r(f)}$ are projections, this is equivalent to $T_{s(e)}T_{r(f)}=0$. Now relation 2 implies $T_{e}T_{f}=T_{e}T_{s(e)}T_{r(f)}T_{f}=0$. Finally, for d notice that if $f=eg$, $d(f)\geq d(g)$. By taking adjoints if needed, we may assume without loss of generality that $d(f)\geq d(g)$. Then either $f\in e\Lambda$ or $e\Lambda\cap f\Lambda=\emptyset$. First suppose $f=eg$, and note $T_{e}^{*}T_{f}=T_{e}^{*}T_{e}T_{g}=T_{s(e)}T_{g}$, which by relation 2 implies $T_{e}^{*}T_{f}=T_{g}$. Now assume $e\Lambda\cap f\Lambda=0$. Then by relation 3, $T_{e}T_{e}^{*}T_{f}T_{f}^{*}=0$. Multiplying by $T_{f}$ on the right and by $T_{e}^{*}$ on the left, we see $T_{e}^{*}T_{e}T_{e}^{*}T_{f}T_{f}^{*}T_{f}=T_{e}^{*}T_{f}=0$, as desired. ∎
Remark 3.20.
In the proof of the previous lemma, we did not use the full power of relation 3, only that $T_{e}T_{e}^{*}T_{f}T_{f}^{*}=0$ when $e\Lambda\cap f\Lambda=\emptyset$. In fact, when $f=eg$, the identities above give $T_{e}T_{e}^{*}T_{f}T_{f}^{*}=T_{e}T_{g}T_{f}^{*}=T_{f}T_{f}^{*}$, so it is enough to use the weaker form of relation 3 by only requiring the range projections of $T_{e}$ and $T_{f}$ to be orthogonal when $e\Lambda\cap f\Lambda=\emptyset$.
An important consequence of parts c and d in the previous result is that every word composed of the generators $\left\{T_{e}:e\in\Lambda\right\}$ of $\mathcal{O}\left(\Lambda\right)$ and their adjoints is of the form $T_{e}T_{f}^{*}$.
Corollary 3.21.
If $\Lambda$ is an ordinal graph and $\mathcal{O}\left(\Lambda\right)$ has generators $\left\{T_{e}:e\in\Lambda\right\}$, then
$\mathcal{O}\left(\Lambda\right)=\overline{\text{$\mathrm{span}$}}\left\{T_{e}T_{f}^{*}:e,f\in\Lambda\right\}$
Proof.
By the construction of $\mathrm{C}^{*}$-algebras which are universal for a set of generators and relations, the span of words in the elements $\left\{T_{e}:e\in\Lambda\right\}\cup\left\{T_{e}^{*}:e\in\Lambda\right\}$ is dense in $\mathcal{O}\left(\Lambda\right)$. As mentioned above, 3.19 part d implies that every such word is equal to an element in $\left\{T_{e}T_{f}^{*}:e,f\in\Lambda\right\}$, and the result follows. ∎
Let $\alpha,\beta\in\mathrm{Ord}$ with $\alpha\leq\beta$ be given. By the following lemma, each instance in $\mathcal{O}\left(\Lambda_{\alpha}\right)$ of relation 4 from 3.18 continues to hold in $\mathcal{O}\left(\Lambda_{\beta}\right)$. Likewise, relations 1-3 in $\mathcal{T}\left(\Lambda_{\alpha}\right)$ continue to hold in $\mathcal{T}\left(\Lambda_{\beta}\right)$, giving us the inductive system in 3.23.
Lemma 3.22.
If $\alpha\leq\beta$ and $F\subseteq v\Lambda_{\alpha}$ is finite and exhaustive for $v$ in $\Lambda_{\alpha}$, then $F$ is exhaustive for $v$ in $\Lambda_{\beta}$.
Proof.
Suppose $f\in v\Lambda_{\beta}\backslash v\Lambda_{\alpha}$. Factor $f$ as $f=gk$ where $\omega^{\alpha}>d(g)\geq\max_{e\in F}d(e)$. Now for some $e\in F$, $e\Lambda_{\alpha}\cap g\Lambda_{\alpha}\not=\emptyset$, so choose $h\in e\Lambda_{\alpha}\cap g\Lambda_{\alpha}$. Then $h=ea=gb$ for $a,b\in\Lambda_{\alpha}$. Let $g=e^{\prime}c$ where $d(e^{\prime})=d(e)$, whence $h=ea=e^{\prime}cb$. By uniqueness of factorization, $e=e^{\prime}$, and $f\in e\Lambda_{\beta}=e\Lambda_{\beta}\cap f\Lambda_{\beta}$. ∎
Proposition 3.23.
For each $\alpha,\beta\in\mathrm{Ord}$ with $\alpha\leq\beta$ there exist *-homomorphisms $\sigma_{\alpha}^{\beta}:\mathcal{T}\left(\Lambda_{\alpha}\right)\rightarrow\mathcal{T}\left(\Lambda_{\beta}\right)$ and $\rho_{\alpha}^{\beta}:\mathcal{O}\left(\Lambda_{\alpha}\right)\rightarrow\mathcal{O}\left(\Lambda_{\beta}\right)$ defined by
$\sigma_{\alpha}^{\beta}\left(T_{e}\right)$ $=S_{e}$
$\rho_{\alpha}^{\beta}\left(T_{e}\right)$ $=S_{e}$
for $e\in\Lambda_{\alpha}$, where $\left\{T_{e}:e\in\Lambda\right\}$ and $\left\{S_{e}:e\in\Lambda\right\}$ are the corresponding generators of $\mathcal{T}\left(\Lambda\right)$ or $\mathcal{O}\left(\Lambda\right)$. If $\alpha\leq\gamma\leq\beta$, $\sigma_{\gamma}^{\beta}\circ\sigma_{\alpha}^{\gamma}=\sigma_{\alpha}^{\beta}$ and $\rho_{\gamma}^{\beta}\circ\rho_{\alpha}^{\gamma}=\rho_{\alpha}^{\beta}$. If $\beta$ is a limit ordinal, then
$\mathcal{O}\left(\Lambda_{\beta}\right)$ $=\overline{\bigcup_{\alpha<\beta}\rho_{\alpha}^{\beta}\left(\mathcal{O}\left(\Lambda_{\alpha}\right)\right)}$
$\mathcal{T}\left(\Lambda_{\beta}\right)$ $=\overline{\bigcup_{\alpha<\beta}\sigma_{\alpha}^{\beta}\left(\mathcal{T}\left(\Lambda_{\alpha}\right)\right)}$
Proof.
The existence of the inductive systems follows from 3.22, the universal properties of $\mathcal{T}\left(\Lambda\right)$ and $\mathcal{O}\left(\Lambda\right)$, and the discussion above. Let $\left\{S_{e}:e\in\Lambda_{\beta}\right\}$ be the generators of $\mathcal{O}\left(\Lambda_{\beta}\right)$. If $\alpha<\beta$ and $\beta$ is a limit ordinal, $\alpha+1<\beta$. Hence if $d\left(e\right)<\beta$, $S_{e}=\rho_{d\left(e\right)+1}^{\beta}\left(V_{e}\right)$ where $V_{e}$ is the generator for $e$ in $\mathcal{O}\left(\Lambda_{d\left(e\right)+1}\right)$. Thus $\cup_{\alpha<\beta}\rho_{\alpha}^{\beta}\left(\mathcal{O}\left(\Lambda_{\alpha}\right)\right)$ contains all finite sums of words in the generators, and hence is dense in $\mathcal{O}\left(\Lambda_{\beta}\right)$. Identical logic applies to $\mathcal{T}\left(\Lambda_{\beta}\right)$. ∎
Next we define the connected components of an ordinal graph. The connected components of an ordinal graph are the finest partition of the morphisms for which composable $f,g\in\Lambda$ belong to the same part. Each connected component of an ordinal graph is also an ordinal graph. We wish to describe the relationship between the Cuntz-Krieger algebra of $\Lambda$ and the Cuntz-Krieger algebras of its connected components.
Definition 3.24.
For an ordinal graph $\Lambda$, define the relation $\preceq$ on $\Lambda$ where $e\preceq f$ if there is $g\in\Lambda$ with $s(g)=r(f)$ and $r(g)=s(e)$. Let $\sim$ be the equivalence relation generated by $\preceq$, that is, $e\sim f$ if for every equivalence relation $\simeq$ on $\Lambda$ with the property that $g\preceq h$ implies $g\simeq h$, $e\simeq f$. Then $F\subseteq\Lambda$ is a connected component of $\Lambda$ if $F$ is an equivalence class for $\sim$. Let $\mathcal{F}\left(\Lambda\right)$ denote the set of connected components of $\Lambda$.
Proposition 3.25.
If $\Lambda$ is an ordinal graph, then each $F\in\mathcal{F}\left(\Lambda\right)$ is an ordinal graph. Moreover, we have the following isomorphism.
$\mathcal{O}\left(\Lambda\right)\cong\bigoplus_{F\in\mathcal{F}\left(\Lambda\right)}^{c_{0}}\mathcal{O}\left(F\right)$
Proof.
Let $F\in\mathcal{F}\left(\Lambda\right)$ and $e,f\in F$ with $s(e)=r(f)$. Then $ef\in F$. Moreover, if $\alpha\leq d(e)$, $r\left(e_{\alpha}\right)=r\left(e\right)$ and $s\left(e^{\alpha}\right)=s\left(e\right)$, so $e_{\alpha},e^{\alpha}\in F$. Thus $F$ is an ordinal graph.
Let $\mathcal{A}=\oplus_{F\in\mathcal{F}}\mathcal{O}\left(F\right)$. We will construct a Cuntz-Krieger $\Lambda$-family in $\mathcal{A}$. For $e\in\Lambda$, denote by $[e]$ the connected component of $\Lambda$ containing $e$. Let $\left\{T_{e}:e\in F\right\}$ be the generators of $\mathcal{O}\left(F\right)$ for each $F\in\mathcal{F}\left(\Lambda\right)$, and let $\left\{V_{e}:e\in\Lambda\right\}$ be the generators for $\mathcal{O}\left(\Lambda\right)$. Then for each $e\in\Lambda$, define $S_{e}\in\mathcal{A}$ by
$\left(S_{e}\right)_{F}=\begin{cases}T_{e}&[e]=F\\ 0&\text{otherwise}\end{cases}$
If $e\in\Lambda$, $[s(e)]=[e]$, so indeed $S_{e}^{*}S_{e}=S_{s(e)}$. Likewise, if $e,f\in\Lambda$ with $s(e)=r(f)$, $[e]=[f]$ so $S_{e}S_{f}=S_{ef}$. If $s(e)\not=r(f)$ and $[e]=[f]$, $S_{e}S_{f}=0$ since $T_{e}T_{f}=0$. And if $[e]\not=[f]$, $S_{e}S_{f}=0$. Finally, if $G\subseteq\Lambda$ is finite and exhaustive for $v\in\Lambda_{0}$, then $G\subseteq[v]$. Thus $\sum_{e\in G}S_{e}S_{e}^{*}=S_{v}$, and $\left\{S_{e}:e\in\Lambda\right\}$ is a Cuntz-Krieger $\Lambda$-family. This gives a surjective *-homomorphism $\mu:\mathcal{O}\left(\Lambda\right)\rightarrow\mathcal{A}$ such that $\mu\left(V_{e}\right)=S_{e}$.
If $F\in\mathcal{F}\left(\Lambda\right)$, $\left\{V_{e}:e\in F\right\}$ is a Cuntz-Krieger $F$-family. Thus there exists $\nu_{F}:\mathcal{O}\left(F\right)\rightarrow\mathcal{O}\left(\Lambda\right)$ where $\nu_{F}\left(T_{e}\right)=V_{e}$. Hence $\left(\mu\circ\nu_{F}\right)_{F}=\text{id}_{\mathcal{O}\left(F\right)}$, and $\nu_{F}$ is injective. For $e\in\Lambda$, $\nu_{F}\left(\mathcal{O}\left(F\right)\right)T_{e}=0$ if $e\not\in F$, so $\left\{\nu_{F}\left(\mathcal{O}\left(F\right)\right):F\in\mathcal{F}\left(\Lambda\right)\right\}$ is a family of pairwise orthogonal ideals which generate $\mathcal{O}\left(\Lambda\right)$. Therefore $\mathcal{O}\left(\Lambda\right)\cong\oplus_{F\in\mathcal{F}\left(\Lambda\right)}\nu_{F}\left(\mathcal{O}\left(F\right)\right)\cong\mathcal{A}$. ∎
4. Exhaustive Sets
Let $\Lambda$ be a fixed ordinal graph. In this section, we wish to understand the exhaustive sets which appear in relation 4 of 3.18 and are used to define $\mathcal{O}\left(\Lambda\right)$. By assuming relation 4 only holds for some finite exhaustive sets, we recover the relation for all other finite exhaustive sets.
Definition 4.1.
For $\alpha\in\mathrm{Ord}$, $v\in\Lambda_{0}$ is an $\alpha$ source if there is no path $e\in v\Lambda$ such that $d\left(e\right)=\omega^{\alpha}$. $v$ is $\alpha$ source-regular if for every $e\in v\Lambda_{\alpha}$ there is $f\in s\left(e\right)\Lambda$ such that $d\left(f\right)=\omega^{\alpha}$.
Proposition 4.2.
If $v\in\Lambda_{0}$ is an $\alpha$ source and $\beta\geq\alpha$, then $v$ is a $\beta$ source.
Proof.
Suppose $v\in\Lambda_{0}$ and $e\in v\Lambda$ with $d\left(e\right)=\omega^{\beta}$. Since $\omega^{\alpha}\leq\omega^{\beta}$, $r\left(e^{\omega^{\alpha}}\right)=r\left(e\right)=v$. Thus $e^{\omega^{\alpha}}\in v\Lambda$. ∎
If $\Lambda$ is a directed graph, we only require that $\sum_{e\in v\Lambda}T_{e}T_{e}^{*}$ if $v$ is row-finite and not a source. The above definition of $\alpha$ source plays the role of sources for a directed graph. In fact, if $\Lambda$ is a directed graph, then $v$ is a source if and only if $v$ is a 00 source. Likewise, the following definition is analogous to row-finiteness for directed graphs.
Definition 4.3.
For $\alpha\in\mathrm{Ord}$, $v\in\Lambda_{0}$ is $\alpha$ row-finite if |$\left|\left\{f\in v\Lambda:d(f)=\omega^{\alpha}\right\}\right|<\infty$. $v$ is $\alpha$ regular if $v$ is $\alpha$ source-regular and $v$ is $\alpha$ row-finite.
Example 4.4.
Let $\Lambda$ be the ordinal graph in Fig. 3.1. Then $v$ is 1 source-regular and 1 row-finite, hence $v$ is 1-regular.
Example 4.5.
Consider an ordinal graph $\Lambda$ generated by $v,e,f,g$ where
  1. (1)
    $v=s\left(e\right)=r\left(e\right)=s\left(f\right)=r\left(f\right)=s\left(g\right)=r\left(g\right)$
  2. (2)
    $efg=g$
  3. (3)
    $d\left(v\right)=0$
  4. (4)
    $d\left(e\right)=d\left(f\right)=1$
  5. (5)
    $d\left(g\right)=\omega$
Then $v$ is 1 source-regular because if $d\left(h\right)<\omega$ with $r\left(h\right)=v$, $s\left(h\right)=r\left(g\right)$ and $d\left(g\right)=\omega$. However, $v$ is not $1$ row-finite because for each $k\in\mathbb{N}$ there exists a distinct path $\underbrace{ee\ldots e}_{k\text{ times}}g$ of length $\omega$. Thus $v$ is not $1$-regular.
$v$$e$$f$$g:efef\ldots;g=efg$
Figure 4.1. $\Lambda$ as in Example 4.5
The previous example demonstrates how $1$-regularity is much more delicate than 00-regularity. The source of $1$-irregularity is that $\left|v\Lambda^{\omega}\right|=\infty$, and this happens even when $\Lambda$ is finitely generated. On the other hand, a finitely generated ordinal graph will never have $\left|v\Lambda^{1}\right|=\infty$, so 00-regularity occurs often. We dedicate most of this section to proving that all finite exhaustive sets are obtained by considering the paths in $v\Lambda^{\omega^{\alpha}}$ for $\alpha$ regular vertices $v$.
Proposition 4.6.
If $v$ is $\alpha$ regular and $\beta<\alpha$, then $v$ is $\beta$ regular.
Proof.
Suppose $v\in\Lambda_{0}$ is $\alpha$ regular. Let $e\in v\Lambda_{\beta}$ be arbitrary. Then $e\in v\Lambda_{\alpha}$, and since $v$ is $\alpha$ regular, there exists $f\in s\left(e\right)\Lambda$ with $d\left(f\right)=\omega^{\alpha}$. Then $f_{\omega^{\beta}}\in s\left(e\right)\Lambda$, hence $v$ is $\beta$ source-regular. Now suppose for each $k\in\mathbb{N}$ there are distinct $g_{k}\in v\Lambda$ with $d\left(g_{k}\right)=\omega^{\beta}$. Since $v$ is $\alpha$ source-regular, for each $k\in\mathbb{N}$ choose $h_{k}\in s\left(g_{k}\right)\Lambda$ with $d\left(h_{k}\right)=\omega^{\alpha}$. Then $d\left(g_{k}h_{k}\right)=d\left(g_{k}\right)+d\left(h_{k}\right)=\omega^{\beta}+\omega^{\alpha}=\omega^{\alpha}$, and by unique factorization,
|$\left|\left\{g_{k}h_{k}:k\in\mathbb{N}\right\}\right|=\infty$
This contradicts $\alpha$ regularity of $v$. ∎
Lemma 4.7.
If $e\Lambda\cap f\Lambda\not=\emptyset$ and $d(e)=d(f)$, then $e=f$.
Proof.
By 3.17, we may assume without loss of generality that $e\in f\Lambda$. Then $e=fg$ for some $g\in\Lambda$. Since $d\left(es\left(e\right)\right)=d\left(e\right)=d\left(f\right)$ and $es\left(e\right)=fg$, unique factorization implies $e=f$. ∎
Since we wish to see that $\alpha$ regular vertices play the role of regular vertices in directed graphs, we show that in the following sense, relation 4 is only applied to vertices which are $\alpha$-regular for some $\alpha$. Later we will see that it suffices to only consider exhaustive sets whose members have length which is a power of $\omega$.
Theorem 4.8.
$v\in\Lambda_{0}$ is $\alpha$ regular if and only if there exists $F\subseteq v\Lambda^{\omega^{\alpha}}$ which is finite and exhaustive for $v$. If $v$ is $\alpha$ regular, then $v\Lambda^{\omega^{\alpha}}$ is the unique such $F\subseteq v\Lambda^{\omega^{\alpha}}$ which is finite and exhaustive.
Proof.
First assume $v$ is $\alpha$ regular, define $F=v\Lambda^{\omega^{\alpha}}$, and suppose $f\in v\Lambda$. Since $v$ is $\alpha$ row-finite, $F$ is finite. Moreover, either $d(f)<\omega^{\alpha}$ or $d(f)\geq\omega^{\alpha}$. In the first case, $v$ is $\alpha$ source-regular, so there exists $g\in\Lambda^{\omega^{\alpha}}$ such that $r(g)=s(f)$. Then $d(fg)=d(f)+d(g)=d(f)+\omega^{\alpha}=\omega^{\alpha}$. Thus $fg\in F$ with $f\Lambda\cap fg\Lambda\not=\emptyset$. In the second case, factor $f$ as $f=f_{\omega^{\alpha}}f^{\omega^{\alpha}}$ where $d\left(f_{\omega^{\alpha}}\right)=\omega^{\alpha}$. Then $f_{\omega^{\alpha}}\in F$ with $f\Lambda\cap f_{\omega^{\alpha}}\Lambda\not=0$, and $F$ is exhaustive.
Now suppose $G\subseteq v\Lambda^{\omega^{\alpha}}$ is finite and exhaustive for $v$. Then if $g\in\Lambda^{\omega^{\alpha}}$ with $r(g)=v$, there is $f\in G$ such that $f\Lambda\cap g\Lambda\not=\emptyset$. Since $d(f)=d(g)$, 4.7 implies $f=g$. Thus $v$ is $\alpha$ row-finite, and $F=G$. If $e\in v\Lambda_{\alpha}$, then choose $p\in G$ such that $e\Lambda\cap p\Lambda\not=\emptyset$. Since $d(e)<d(p)$, 3.17 implies $p\in e\Lambda$. Let $q\in\Lambda$ with $p=eq$. Then $d(p)=\omega^{\alpha}=d(e)+d(q)$, so $d(q)=\omega^{\alpha}$. Thus $v$ is $\alpha$ source-regular, hence regular. ∎
Definition 4.9.
If $F\subseteq v\Lambda$ is exhaustive for $v\in\Lambda_{0}$, we say $F$ is minimal if whenever $e,f\in F$ and $e\Lambda\cap f\Lambda\not=\emptyset$, $e=f$.
If we start with a finite exhaustive set, we may consider what happens when we continually remove one path in every pair $e,f\in\Lambda$ with $e\Lambda\cap f\Lambda\not=\emptyset$. Eventually we get a minimal exhaustive set, and for these sets, the join of the range projections is the sum of the range projections, as the following two results demonstrate:
Lemma 4.10.
If $e\Lambda\cap f\Lambda\not=\emptyset$ and $e\in g\Lambda$, then $f\Lambda\cap g\Lambda\not=\emptyset$.
Proof.
Suppose $e=gh$, and first assume $f\in e\Lambda$. Then $f\in gh\Lambda\subseteq g\Lambda$, so $f\in f\Lambda\cap g\Lambda$. Now assume $e\in f\Lambda$. Then there is $k\in\Lambda$ such that $e=gh=fk$. Either $d(g)<d(f)$ or $d(g)\geq d(f)$. In the first case, factor $f$ as $f=pq$ where $d(p)=d(g)$. Then by unique factorization, $p=g$ and $f\in g\Lambda$. If $d(g)\geq d(f)$, then factor $g$ as $g=pq$ where $d(p)=d(f)$. Unique factorization similarly implies $p=f$, so $g\in f\Lambda$. In any case, $g\Lambda\cap f\Lambda\not=\emptyset$. ∎
Lemma 4.11.
If $F\subseteq v\Lambda$ is finite, minimal, and exhaustive for $v$ and $\left\{T_{e}:e\in\Lambda\right\}$ is a Toeplitz $\Lambda$-family, then
$\bigvee_{e\in F}T_{e}T_{e}^{*}=\sum_{e\in F}T_{e}T_{e}^{*}$
Proof.
We will induct on subsets of $F$. Suppose $G\subsetneq F$ and
$\bigvee_{e\in G}T_{e}T_{e}^{*}=\sum_{e\in G}T_{e}T_{e}^{*}$
Let $f\in F\backslash G$. Since $F$ is minimal, we have for all $e\in G$
$T_{e}T_{e}^{*}T_{f}T_{f}^{*}=0$
Thus
$\bigvee_{e\in G\cup\{f\}}T_{e}T_{e}^{*}$ $=\left(\bigvee_{e\in G}T_{e}T_{e}^{*}\right)\vee T_{f}T_{f}^{*}=\left(\sum_{e\in G}T_{e}T_{e}^{*}\right)\vee T_{f}T_{f}^{*}$
$=\sum_{e\in G}T_{e}T_{e}^{*}+T_{f}T_{f}^{*}-\sum_{e\in G}T_{e}T_{e}^{*}T_{f}T_{f}^{*}$
$=\sum_{e\in G}T_{e}T_{e}^{*}+T_{f}T_{f}^{*}=\sum_{e\in G\cup\{f\}}T_{e}T_{e}^{*}$
Now we show that having relation 4 for all finite minimal exhaustive sets implies relation 4 holds for all finite exhaustive sets.
Proposition 4.12.
Suppose $\left\{T_{e}:e\in\Lambda\right\}$ is a Toeplitz $\Lambda$-family and for every $v\in\Lambda_{0}$ and $F\subseteq v\Lambda_{0}$ which is finite, minimal, and exhaustive for $v$,
$\sum_{f\in F}T_{f}T_{f}^{*}=T_{v}$
Then $\left\{T_{e}:e\in\Lambda\right\}$ is a Cuntz-Krieger $\Lambda$-family.
Proof.
Suppose $F\subseteq v\Lambda$ is finite and exhaustive for $v$. If $F$ is minimal, then indeed by 4.11,
$\bigvee_{g\in F}T_{g}T_{g}^{*}=\sum_{g\in F}T_{g}T_{g}^{*}=T_{v}$
Otherwise, choose $e,f\in F$ such that $f\in e\Lambda$. Then
$\bigvee_{g\in F}T_{g}T_{g}^{*}$ $=\left(\bigvee_{g\in F\backslash\{e,f\}}T_{g}T_{g}^{*}\right)\vee T_{e}T_{e}^{*}\vee T_{f}T_{f}^{*}$
$=\left(\bigvee_{g\in F\backslash\{e,f\}}T_{g}T_{g}^{*}\right)\vee\left(T_{e}T_{e}^{*}+T_{f}T_{f}^{*}-T_{e}T_{e}^{*}T_{f}T_{f}^{*}\right)$
$=\left(\bigvee_{g\in F\backslash\{e,f\}}T_{g}T_{g}^{*}\right)\vee\left(T_{e}T_{e}^{*}+T_{f}T_{f}^{*}-T_{f}T_{f}^{*}\right)=\bigvee_{g\in F\backslash\{f\}}T_{g}T_{g}^{*}$
By 4.10, if $h\Lambda\cap f\Lambda\not=\emptyset$, then $h\Lambda\cap e\Lambda\not=\emptyset$. Thus $F\backslash\left\{f\right\}$ is exhaustive. If $F\backslash\{f\}$ is not minimal, then the process above may be repeated. Since $\left|F\backslash\left\{f\right\}\right|<\left|F\right|$ and $\left|F\right|$ is finite, the process eventually terminates with a finite exhaustive minimal $F^{\prime}$ such that
$\bigvee_{g\in F}T_{g}T_{g}^{*}=\bigvee_{g\in F^{\prime}}T_{g}T_{g}^{*}=\sum_{g\in F^{\prime}}T_{g}T_{g}^{*}=T_{v}$
Once we have a finite minimal exhaustive set $F$ for $v$, we can further reduce relation 4 by splitting $F$ into two smaller finite minimal exhaustive sets $F_{e}$ and $F^{e}$. If relation 4 holds for $F_{e}$ and $F^{e}$, then relation 4 holds for $F$, as shown in the next two results.
Proposition 4.13.
Suppose $F\subseteq v\Lambda$ is minimal and exhaustive for $v$. Let $e\in v\Lambda$ and suppose $ef\in F$ for some $f\in\Lambda$. Define
$F_{e}$ $=\left(F\backslash e\Lambda\right)\cup\{e\}$
$F^{e}$ =$=\left\{g\in\Lambda:s\left(e\right)=r\left(g\right)\text{ and }eg\in F\right\}$
Then $F_{e}$ and $F^{e}$ are minimal. $F_{e}$ is exhaustive for $v$ and $F^{e}$ is exhaustive for $s(e)$.
Proof.
Let $k\in v\Lambda$ be given. Choose $p\in F$ such that $p\Lambda\cap k\Lambda\not=\emptyset$. If $p\not\in e\Lambda$, then $p\in F_{e}$. Otherwise by 4.10, $e\Lambda\cap k\Lambda\not=\emptyset$. Therefore, $F_{e}$ is exhaustive. We wish to prove $F_{e}$ is minimal. To do so, assume $p,q\in F_{e}$ and $p\Lambda\cap q\Lambda\not=\emptyset$. If $p\in F\backslash e\Lambda$ and $q=e$, then by 3.17, $e\in p\Lambda$. This means $ef\in p\Lambda$, and since $F$ is minimal, $ef=p$, contradicting $p\not\in e\Lambda$. Thus we may assume without loss of generality that either $\left\{p,q\right\}\subseteq F\backslash e\Lambda$ or $p=q=e$. In the first case, minimality of $F$ implies $p=q$, so in any case we have $p=q$. Thus $F_{e}$ is minimal.
Now, let $k\in s(e)\Lambda$ be given. Choose $p\in F$ such that $p\Lambda\cap ek\Lambda\not=\emptyset$. If $p\in ek\Lambda$, choose $q\in\Lambda$ with $p=ekq$. Then $kq\in F^{e}$ and $kq\in k\Lambda\cap kq\Lambda$. On the other hand, if $ek\in p\Lambda$, choose $q\in\Lambda$ with $ek=pq$. Then $ek\in e\Lambda\cap p\Lambda$, so by 3.17, either $p\in e\Lambda$ or $e\in p\Lambda$. In the first case, choose $j\in\Lambda$ such that $p=ej$. Since $ek=ejq$, $k=jq$ and $k\in j\Lambda$. Moreover, $j\in F^{e}$ since $p\in F$. In the second case, $e\in p\Lambda$, hence $ef\in p\Lambda$. Minimality of $F$ implies $ef=p$, hence $ek=efq$, and $k=fq\in f\Lambda$. As $f\in F^{e}$, this implies $F^{e}$ is exhaustive for $s(e)$.
Finally, suppose $g,h\in F^{e}$ with $g\Lambda\cap h\Lambda\not=\emptyset$. Then $eg\Lambda\cap eh\Lambda\not=\emptyset$, so by minimality of $F$, $eg=eh$. Then $g=h$, and $F^{e}$ is minimal. ∎
Proposition 4.14.
Suppose $\left\{T_{e}:e\in\Lambda\right\}$ is a Toeplitz $\Lambda$-family. Assume $v\in\Lambda_{0}$, $e\in v\Lambda$, $F\subseteq v\Lambda$ with $F$ finite, minimal, and exhaustive for $v$, and $F\cap e\Lambda\not=\emptyset$. If
$\sum_{f\in F_{e}}T_{f}T_{f}^{*}$ $=T_{v}$
$\sum_{f\in F^{e}}T_{f}T_{f}^{*}$ $=T_{s(e)}$
then
$\sum_{f\in F}T_{f}T_{f}^{*}=T_{v}$
Proof.
We compute
$\sum_{f\in F}T_{f}T_{f}^{*}$ $=\sum_{f\in F\cap e\Lambda}T_{f}T_{f}^{*}+\sum_{f\in F\backslash e\Lambda}T_{f}T_{f}^{*}$
$=\sum_{g\in F^{e}}T_{eg}T_{eg}^{*}+\sum_{f\in F_{e}}T_{f}T_{f}^{*}-T_{e}T_{e}^{*}$
$=T_{e}\left(\sum_{g\in F^{e}}T_{g}T_{g}^{*}\right)T_{e}^{*}-T_{e}T_{e}^{*}+\sum_{f\in F_{e}}T_{f}T_{f}^{*}$
$=T_{e}T_{e}^{*}-T_{e}T_{e}^{*}+\sum_{f\in F_{e}}T_{f}T_{f}^{*}=\sum_{f\in F_{e}}T_{f}T_{f}^{*}=T_{v}$
Lemma 4.15.
If $\alpha\in\mathrm{Ord}$ and $F\subseteq v\Lambda^{\alpha}$ is exhaustive for $v$, then $F$ is minimal.
Proof.
Suppose without loss of generality that $e,f\in F$ and $f\in e\Lambda$. Write $f=eg$ for some $g\in\Lambda$. Then
$d\left(f\right)=\alpha=d\left(e\right)+d\left(g\right)=\alpha+d\left(g\right)$
Hence $d(g)=0$, and $eg=e$. ∎
This suggests that we consider only exhaustive sets whose members have a fixed length. In fact, knowing relation 4 holds for these sets implies it holds for all finite exhaustive sets.
Proposition 4.16.
Suppose $\left\{T_{e}:e\in\Lambda\right\}$ is a Toeplitz $\Lambda$-family such that for all $v\in\Lambda_{0}$, $\alpha\in\mathrm{Ord}$, and finite $F\subseteq v\Lambda^{\alpha}$ which are exhaustive for $v$,
$\sum_{e\in F}T_{e}T_{e}^{*}=T_{v}$
Then $\left\{T_{e}:e\in\Lambda\right\}$ is a Cuntz-Krieger $\Lambda$-family.
Proof.
By 4.12, it suffices to let $F$ be an arbitrary finite, minimal, exhaustive set for $v$ and induct on $|F|$ to prove that $\sum_{f\in F}T_{f}T_{f}^{*}=T_{v}$. Towards that end, note that if $|F|=1$ and $f\in F$, then by hypothesis $T_{f}T_{f}^{*}=T_{v}$. Now suppose that $G$ is minimal and exhaustive for $v$, $|G|=n$, and that for all $u\in\Lambda_{0}$ with minimal and exhaustive $F$ for $u$ such that $|F|<n$, $\sum_{f\in F}T_{f}T_{f}^{*}=T_{u}$. Let $m=\min_{g\in G}d\left(g\right)$, and label $G$ as $G=\left\{g_{1},\ldots g_{n}\right\}$ where $d\left(g_{k}\right)\leq d\left(g_{k+1}\right)$. For each $1\leq k\leq n$, factor $g_{k}$ as $g_{k}=p_{k}q_{k}$, where $d\left(p_{k}\right)=m$. Define $G_{0}=G$, $G_{1}=G_{p_{1}}$, and note that since $G$ is minimal and $g_{1}=p_{1}$, $G_{1}=G$. For $k\geq 1$, define $G_{k}=\left(G_{k-1}\right)_{p_{k}}$, $G^{k}=\left(G_{k-1}\right)^{p_{k}}$. Then $G_{n}=\left\{p_{1},\ldots p_{n}\right\}\subseteq\Lambda^{m}$ is minimal and exhaustive for $v$, so by hypothesis
$\sum_{k=1}^{n}T_{p_{k}}T_{p_{k}}^{*}=T_{v}$
Moreover, if $1\leq k\leq n$, $G^{k}$ is minimal and exhaustive for $s\left(p_{k}\right)$. If $\left|G^{k}\right|<n$ for $1\leq k\leq n$, then by the induction hypothesis,
$\sum_{f\in G^{k}}T_{f}T_{f}^{*}=T_{s\left(p_{k}\right)}$
By repeatedly applying 4.14, we see $G_{n-1},G_{n-2},\ldots,G_{1}$ satisfy relation 4. That is, for $1\leq k\leq n$,
$\sum_{f\in G_{k}}T_{f}T_{f}^{*}=T_{v}$
Since $G_{1}=G$, we have the desired identity.
All that remains is the case when $\left|G^{k}\right|=n$ for some $1<k\leq n$. First we prove that $g_{1}\in G_{k}$ for $1\leq k\leq n$. Since $G_{1}=G_{p_{1}}$, $g_{1}=p_{1}$, and $g_{1}\in G$, we have $g_{1}\in G_{1}$. Moreover if $g_{1}\in G_{k}\backslash G_{k+1}$, then $g_{1}=p_{1}\in p_{k+1}\Lambda$. As $d\left(p_{k+1}\right)\geq d\left(p_{1}\right)$, this implies $p_{1}=p_{k+1}$. Then $g_{k+1}=p_{k+1}q_{k+1}=g_{1}q_{k+1}$, and because $G$ is minimal, $g_{k+1}=g_{1}$. This contradicts the assumption that $\left|G\right|=n$, therefore we must have $g_{1}\in G_{k}$ for all $1\leq k\leq n$. Suppose for some $1<k\leq n$, $\left|G^{k}\right|=n$. Then since $g_{1}\in G_{k-1}$, $p_{1}\in p_{k}\Lambda$. Since $d\left(p_{1}\right)\leq d\left(p_{k}\right)$, this implies $g_{1}=p_{k}$. Hence $g_{k}\in g_{1}\Lambda$, and by minimality of $G$, $g_{1}=g_{k}$. Once again, this contradicts $\left|G\right|=n$, and hence $\left|G^{k}\right|<n$ for $1<k\leq n$. ∎
Finally, we may instead only consider finite and exhaustive sets whose members have lengths which are a fixed power of $\omega$.
Theorem 4.17.
Suppose $\left\{T_{e}:e\in\Lambda\right\}$ is a Toeplitz $\Lambda$-family such that for all $v\in\Lambda_{0}$, $\alpha\in\mathrm{Ord}$, and $F\subseteq v\Lambda^{\omega^{\alpha}}$ finite and exhaustive for $v$,
$\sum_{e\in F}T_{e}T_{e}^{*}=T_{v}$
Then $\left\{T_{e}:e\in\Lambda\right\}$ is a Cuntz-Krieger $\Lambda$-family.
Proof.
We will prove by transfinite induction that for all $v\in\Lambda_{0}$ and $\beta\in\mathrm{Ord}$, if $G\subseteq v\Lambda^{\beta}$ is finite and exhaustive, then
$\sum_{e\in G}T_{e}T_{e}^{*}=T_{v}$
Notice that if $\beta=0$ and $G\subseteq v\Lambda^{\beta}$ is finite and exhaustive, then $G=\{v\}$, so the statement is true. Now suppose the statement is true for all $\gamma<\beta$, and let $G\subseteq v\Lambda^{\beta}$ be finite and exhaustive. First, choose $\alpha\in\mathrm{Ord}$ such that $\omega^{\alpha}\leq\beta<\omega^{\alpha+1}$, and note that 2.1 implies $-\omega^{\alpha}+\beta<\beta$. Label the elements of $G$ as $G=\left\{g_{1},\ldots g_{n}\right\}$. For each $1\leq k\leq n$, factor $g_{k}$ as $g_{k}=p_{k}q_{k}$ where $d\left(p_{k}\right)=\omega^{\alpha}$. Then $d\left(g_{k}\right)=\beta=d\left(p_{k}\right)+d\left(q_{k}\right)=\omega^{\alpha}+d\left(q_{k}\right)$, hence $d\left(q_{k}\right)=-\omega^{\alpha}+\beta$. Define $G_{0}=G$, and for $1\leq k\leq n$, $G_{k}=\left(G_{k-1}\right)_{p_{k}}$ and $G^{k}=\left(G_{k-1}\right)^{p_{k}}$. Then $G_{n}=\left\{p_{1},\ldots p_{n}\right\}$, and $G^{k}\subseteq s\left(p_{k}\right)\Lambda^{-\omega^{\alpha}+\beta}$. By hypothesis, $\sum_{e\in G_{n}}T_{e}T_{e}^{*}=T_{v}$. Moreover, since $-\omega^{\alpha}+\beta<\beta$, the inductive hypothesis implies that $\sum_{e\in G^{k}}T_{e}T_{e}^{*}=T_{s\left(p_{k}\right)}$ for each $k$. Applying 4.14 repeatedly, we see
$\sum_{e\in G_{k}}T_{e}T_{e}^{*}=T_{v}$
for each $k$, and in particular, for $k=0$. ∎
Corollary 4.18.
$\mathcal{O}\left(\Lambda\right)$ is the universal $C^{*}$-algebra generated by a family of operators $\left\{T_{e}:e\in\Lambda\right\}$ satisfying the following relations:
  1. (1)
    $T_{e}^{*}T_{e}=T_{s(e)}$
  2. (2)
    $T_{e}T_{f}=\begin{cases}T_{ef}&s(e)=r(f)\\ 0&\text{otherwise}\end{cases}$
  3. (3)
    $T_{e}T_{e}^{*}T_{f}T_{f}^{*}=T_{f}T_{f}^{*}$ if $f\in e\Lambda$
  4. (4)
    $T_{v}=\sum_{e\in v\Lambda^{\omega^{\alpha}}}T_{e}T_{e}^{*}$ if $v$ is $\alpha$ regular
Proof.
By 4.8, if $F\subseteq v\Lambda^{\omega^{\alpha}}$ is finite and exhaustive for $v$, then $F=v\Lambda^{\omega^{\alpha}}$ and $v$ is $\alpha$ regular. Thus by 4.17, the $\mathrm{C}^{*}$-algebra generated by the relations above is a Cuntz-Krieger $\Lambda$-family. Moreover, for every $\alpha$ regular vertex $v$, $v\Lambda^{\omega^{\alpha}}$ is finite and exhaustive, and thus the representation of $\mathcal{O}\left(\Lambda\right)$ is an isomorphism. ∎
Frequently, it is convenient to regard $\mathcal{O}\left(\Lambda\right)$ as a $\mathrm{C}^{*}$-algebra generated by the paths in $\Lambda^{\omega^{\alpha}}$ instead of all paths. In the following results, we give generators and relations for this context.
Lemma 4.19.
Each $e\in\Lambda$ with $d(e)>0$ can be written uniquely as $e=f_{1}f_{2}\ldots f_{n}$ where for each $k$, $f_{k}\in\Lambda^{\omega^{\beta_{k}}}$ and $\beta_{k}\geq\beta_{k+1}$.
Proof.
For existence, let $e\in\Lambda$ be given, and write $d(e)=\sum_{k=1}^{n}\omega^{\beta_{k}}$ in Cantor normal form using 2.1. We will induct on $n$. Note that if $n=1$, $d(e)=\omega^{\beta_{1}}$. Setting $n=1$ and $f_{1}=e$, we get the desired factorization, and this factorization is unique. Now, suppose the lemma is true for factorizations of length $m<n$. Then
$d(e)=d\left(e_{\omega^{\beta_{1}}}e^{\omega^{\beta_{1}}}\right)=\omega^{\beta_{1}}+\sum_{k=2}^{n}\omega^{\beta_{k}}$
By the induction hypothesis, $e^{\omega^{\beta_{1}}}$ can be uniquely factored as $f_{2}\ldots f_{n}$ where $d\left(f_{k}\right)=\omega^{\beta_{k}}$ for $n\geq k\geq 2$. Setting $f_{1}=e_{\omega^{\beta_{1}}}$, we get a factorization $e=f_{1}f_{2}\ldots f_{n}$, and by 3.1, this factorization is unique. ∎
Proposition 4.20.
$\mathcal{O}\left(\Lambda\right)$ is the universal $C^{*}$-algebra generated by partial isometries
$\left\{S_{v}:v\in\Lambda_{0}\right\}\cup\bigcup_{\alpha\in\mathrm{Ord}}\left\{S_{e}:e\in\Lambda^{\omega^{\alpha}}\right\}$
satisfying the following relations:
  1. (1)
    $S_{e}^{*}S_{e}=S_{s(e)}$
  2. (2)
    $S_{e}S_{f}=S_{ef}$ if $s\left(e\right)=r\left(f\right)$ and $d\left(e\right)<d\left(f\right)$
  3. (3)
    $S_{e}^{*}S_{f}=0$ if $e\Lambda\cap f\Lambda=\emptyset$
  4. (4)
    $S_{v}=\sum_{e\in v\Lambda^{\omega^{\alpha}}}S_{e}S_{e}^{*}$ if $v$ is $\alpha$-regular
Remark 4.21.
Relation 2 makes sense because if $\beta<\alpha$ with $e\in\Lambda^{\omega^{\beta}}$, $f\in\Lambda^{\omega^{\alpha}}$, then $d\left(ef\right)=d\left(e\right)+d\left(f\right)=\omega^{\beta}+\omega^{\alpha}=\omega^{\alpha}$.
Proof.
Let $\mathcal{A}$ be the $\mathrm{C}^{*}$-algebra which is universal for the generators and relations above. By 4.18, there exists a *-homomorphism $\eta:\mathcal{A}\rightarrow\mathcal{O}\left(\Lambda\right)$ with $\eta\left(S_{e}\right)=T_{e}$ for $e\in\Lambda_{0}\cup\bigcup_{\alpha\in\mathrm{Ord}}\Lambda^{\omega^{\alpha}}$. We will construct an inverse $\mu:\mathcal{O}\left(\Lambda\right)\rightarrow\mathcal{A}$, which will finish the proof. We wish to do so by identifying a family $\mathcal{U}=\left\{U_{e}:e\in\Lambda\right\}$ of operators which satisfy the relations of 4.18. $\mu$ is an inverse for $\eta$ if and only if $\mu\left(T_{e}\right)=S_{e}$ for $e\in\Lambda_{0}\cup\bigcup_{\alpha\in\mathrm{Ord}}\Lambda^{\omega^{\alpha}}$. Thus for such $e$, we must define $U_{e}=S_{e}$. By 4.19 and relation 2 in 4.18, this determines $\mathcal{U}$, for if $e=f_{1}f_{2}\ldots f_{n}$ with $f_{k}\in\Lambda^{\omega^{\beta_{k}}}$, $\beta_{k}\geq\beta_{k+1}$, then
$U_{e}=U_{f_{1}}U_{f_{2}}\ldots U_{f_{n}}=S_{f_{1}}S_{f_{2}}\ldots S_{f_{n}}$
Suppose $g=h_{1}h_{2}\ldots h_{m}$ with $h_{k}\in\Lambda^{\omega^{\alpha_{k}}}$, $\alpha_{k}\geq\alpha_{k+1}$, and $s(g)=r(e)$. Choose the smallest $r$ with $1\leq r\leq m$ such that $\alpha_{r}<\beta_{1}$. With $p=h_{r}h_{r+1}\ldots h_{m}f_{1}$, $d(p)=\omega^{\beta_{1}}$ and
$U_{g}U_{e}$ $=U_{h_{1}}\ldots U_{h_{r}}U_{h_{r+1}}\ldots U_{h_{m}}U_{f_{1}}\ldots U_{f_{n}}$
$=U_{h_{1}}\ldots U_{h_{r-1}}U_{p}U_{f_{2}}\ldots U_{f_{n}}$
Then $\alpha_{1}\geq\alpha_{2}\geq\ldots\geq\alpha_{r-1}\geq\beta_{1}\geq\ldots\geq\beta_{n}$, so by uniqueness in 4.19, $U_{ge}=U_{g}U_{e}$. It remains to see that $\mathcal{U}$ satisfies the other relations of 4.18. We may verify relation 1 with the following calculation:
$U_{e}^{*}U_{e}$ $=U_{f_{n}}^{*}U_{f_{n-1}}^{*}\ldots U_{f_{1}}^{*}U_{f_{1}}\ldots U_{f_{n}}$
$=U_{f_{n}}^{*}\ldots U_{f_{2}}^{*}U_{s\left(f_{1}\right)}U_{f_{2}}\ldots U_{f_{n}}$
$=U_{f_{n}}^{*}\ldots U_{f_{2}}^{*}U_{f_{2}}\ldots U_{f_{n}}$
$\ldots$
$=U_{f_{n}}^{*}U_{f_{n}}=U_{s\left(f_{n}\right)}=U_{s(e)}$
For relation 3, suppose without loss of generality $d(e)\leq d(g)$ and $e\Lambda\cap g\Lambda=\emptyset$. If for $1\leq r\leq n$ and $\beta=\sum_{k=1}^{r}\omega^{\beta_{k}}$, $e_{\beta}=g_{\beta}$, then
$U_{e}^{*}U_{g}$ $=U_{e^{\beta}}^{*}U_{e_{\beta}}^{*}U_{g_{\beta}}U_{g^{\beta}}$
$=U_{e^{\beta}}^{*}U_{s\left(e_{\beta}\right)}U_{g^{\beta}}=U_{e^{\beta}}^{*}U_{g^{\beta}}$
Thus it suffices to show $U_{e^{\beta}}^{*}U_{g^{\beta}}=0$. By applying relation 1, we may also assume without loss of generality that $e_{\omega^{\beta_{1}}}\not=g_{\omega^{\beta_{1}}}$. Then 4.7 implies $e_{\omega^{\beta_{1}}}\Lambda\cap g_{\omega^{\beta_{1}}}\Lambda=\emptyset$, so by relation 3 of $\mathcal{A}$,
$U_{e}^{*}U_{g}=U_{e^{\omega^{\beta_{1}}}}^{*}U_{e_{\omega^{\beta_{1}}}}^{*}U_{g_{\omega^{\beta_{1}}}}U_{g^{\omega^{\beta_{1}}}}=0$
Finally, relation 4 of 4.18 is identical to relation 4 for $\mathcal{A}$. ∎
5. Infinite Paths
For an ordinal graph $\Lambda$, we wish to define the infinite path space on which we represent $\mathcal{O}\left(\Lambda\right)$. This will prove that the generators of $\mathcal{O}\left(\Lambda\right)$ are non-zero. In particular, 7.15 is not useful if any of the vertex projections are zero, but 5.8 proves this never happens.
Definition 5.1.
For an ordinal graph $\Lambda$, define $\Lambda^{*}=\left\{f\in\prod_{\beta<\alpha}\Lambda^{\beta}:\alpha\in\mathrm{Ord},f\left(\beta\right)\in f\left(\gamma\right)\Lambda\text{ if }\gamma\leq\beta\right\}$. For $f\in\Lambda^{*}\cap\prod_{\beta<\alpha}\Lambda^{\beta}$, we define $L\left(f\right)=\alpha$ and $r\left(f\right)=f\left(0\right)$.
If $\Lambda$ is a directed graph, then $\Lambda^{*}$ is the set of finite and infinite paths. For $e\in\Lambda$, we may define $f\in\Lambda^{*}\cap\prod_{\beta<d(e)+1}\Lambda^{\beta}$ by setting $f\left(\beta\right)=e_{\beta}$. Indeed, if $\gamma\leq\beta$, $f\left(\beta\right)=e_{\beta}=\left(e_{\beta}\right)_{\gamma}\left(e_{\beta}\right)^{\gamma}=e_{\gamma}\left(e_{\beta}\right)^{\gamma}\in e_{\gamma}\Lambda=f\left(\gamma\right)\Lambda$. In this way, we regard $\Lambda$ as a subset of $\Lambda^{*}$. Then for $e\in\Lambda$, $L\left(e\right)=d(e)+1$.
$\Lambda^{*}$ is partially ordered, where $e\leq f$ if $L(e)\leq L(f)$ and for $\beta<L(e)$, $e\left(\beta\right)=f\left(\beta\right)$. For a directed graph, the maximal elements of $\Lambda^{*}$ are the infinite paths and the finite paths $e$ for which $s(e)$ is a source.
Definition 5.2 (cf. [8, Definition 2.1]).
$\Lambda^{\infty}$ is the set of maximal elements of $\Lambda^{*}$.
Proposition 5.3.
For every $e\in\Lambda$ there exists $f\in\Lambda^{\infty}$ such that $L\left(f\right)>d\left(e\right)$ and $f\left(d\left(e\right)\right)=e$.
Proof.
Let $e\in\Lambda$ be arbitrary. Define $\Omega=\left\{g\in\Lambda^{*}:L\left(g\right)>d\left(e\right)\text{ and }g\left(d\left(e\right)\right)=e\right\}$. If $f$ is a maximal element of $\Omega$ and $g\in\Lambda^{*}$ with $g\geq f$, then $g\in\Omega$. Thus every maximal element of $\Omega$ belongs to $\Lambda^{\infty}$. Hence it suffices to show $\Omega$ has a maximal element. By Zorn’s lemma, we only need to show that every totally ordered subset of $\Omega$ has an upper bound. Let $C\subseteq\Omega$ be a totally ordered subset of $\Omega$, and define $\alpha=\sup_{h\in C}L(h)$. We also define $f\in\Lambda^{*}\cap\prod_{\beta<\alpha}\Lambda^{\beta}$ where for $\beta<\alpha$, $h\in C$ with $L\left(h\right)>\beta$, $f\left(\beta\right)=h\left(\beta\right)$. By construction, for every $\beta<\alpha$ there exists $h\in C$ such that $L\left(h\right)>\beta$. Moreover, $h\left(\beta\right)$ is independent of the choice of $h$ since $C$ is totally ordered, so $f$ is well-defined. Indeed, we also constructed $f$ to be an upper bound for $C$. ∎
For $e\in\Lambda$ and $f\in\Lambda^{*}$ with $s(e)=r(f)$, we wish to define a composition $ef\in\Lambda^{*}$ which is compatible with composition in $\Lambda$. We achieve this with the following definition and result.
Definition 5.4.
If $e\in\Lambda$ and $f\in\Lambda^{*}$ with $s(e)=r(f)$, define $ef\in\prod_{\beta<d(e)+L(f)}\Lambda^{\beta}$ by
$\left(ef\right)\left(\beta\right)=\begin{cases}e_{\beta}&\beta\leq d\left(e\right)\\ ef\left(-d\left(e\right)+\beta\right)&\beta>d\left(e\right)\end{cases}$
By construction, $ef\in\Lambda^{*}$ if $e\in\Lambda$ and $f\in\Lambda^{*}$ with $s\left(e\right)=r\left(f\right)$. Note also that if $f,g\in\Lambda^{*}$, $e\in\Lambda$, $s(e)=r(f)=r(g)$, and $f\leq g$, then $ef\leq eg$. Moreover, for $f\in\Lambda^{*}$ and $\beta<L\left(f\right)$, there is $g\in\Lambda^{*}$ with $f=f\left(\beta\right)g$, where $g\in\prod_{\gamma<-\beta+L\left(f\right)}\Lambda^{\gamma}$ is defined by
$g\left(\gamma\right)=f\left(\beta+\gamma\right)$
Lemma 5.5.
If $f\in\Lambda^{*}$ and $e\in\Lambda r(f)$, then $ef\in\Lambda^{\infty}$ if and only if $f\in\Lambda^{\infty}$.
Proof.
Suppose $g\in\Lambda^{*}$, $f\in\Lambda^{\infty}$, and $ef\leq g$. Then $L(g)\geq L\left(ef\right)=d(e)+L\left(f\right)\geq d(e)$. Thus $g\left(d\left(e\right)\right)=\left(ef\right)\left(d\left(e\right)\right)=e$, and $g=eh$ where $L\left(h\right)=-d\left(e\right)+L\left(g\right)$ and $h\left(\beta\right)=g\left(d\left(e\right)+\beta\right)$. Hence $ef\leq eh$, and $f\leq h$. Since $f\in\Lambda^{\infty}$, $h=f$ and $g=ef$.
For the other direction, suppose $ef\in\Lambda^{\infty}$ and $g\in\Lambda^{*}$ such that $f\leq g$. Then $r(f)=r(g)=s(e)$, and $ef\leq eg$. Since $ef\in\Lambda^{\infty}$, $ef=eg$, and hence $f=g$. ∎
Lemma 5.6.
If $e,f\in\Lambda$ with $s(e)=r(f)$ and $g\in\Lambda^{*}$ with $s(f)=r(g)$, $\left(ef\right)g=e\left(fg\right)$.
Proof.
First we verify $L\left(\left(ef\right)g\right)=d\left(ef\right)+L\left(g\right)=d\left(e\right)+d\left(f\right)+L\left(g\right)=d\left(e\right)+L\left(fg\right)=L\left(e\left(fg\right)\right)$. Now, let $\beta<L\left(\left(ef\right)g\right)$ be given. If $\beta\leq d\left(e\right)$,
$\left(\left(ef\right)g\right)\left(\beta\right)=\left(ef\right)_{\beta}=e_{\beta}=\left(e\left(fg\right)\right)\left(\beta\right)$
If $d\left(e\right)\leq\beta\leq d\left(ef\right)=d\left(e\right)+d\left(f\right)$, $-d\left(e\right)+\beta\leq d\left(f\right)$ and
$\left(\left(ef\right)g\right)\left(\beta\right)=\left(ef\right)_{\beta}=ef_{-d\left(e\right)+\beta}=e\left(fg\right)\left(-d\left(e\right)+\beta\right)=\left(e\left(fg\right)\right)\left(\beta\right)$
Finally, if $d\left(ef\right)\leq\beta$, then
$\left(\left(ef\right)g\right)\left(\beta\right)$ $=\left(ef\right)g\left(-d\left(ef\right)+\beta\right)=\left(ef\right)g\left(-d(f)-d(e)+\beta\right)$
$=e\left(fg\right)\left(-d\left(e\right)+\beta\right)=\left(e\left(fg\right)\right)\left(\beta\right)$
Lemma 5.7.
If $v\in\Lambda_{0}$ is $\alpha$-regular and $f\in\Lambda^{\infty}$ with $r(f)=v$, then $L\left(f\right)>\omega^{\alpha}$.
Proof.
Since $v$ is $\alpha$-regular, $\left|v\Lambda^{\omega^{\alpha}}\right|<\infty$. Thus if $L\left(f\right)\leq\omega^{\alpha}$, we may define a function $g:\left[0,L\left(f\right)\right)\rightarrow\mathbb{N}$ where $g\left(\beta\right)=\left|f\left(\beta\right)\Lambda^{\omega^{\alpha}}\right|$. Note that $g$ is decreasing because if $\gamma\leq\beta$,
$f\left(\beta\right)\Lambda^{\omega^{\alpha}}=f\left(\gamma\right)f\left(\beta\right)^{\gamma}\Lambda^{\omega^{\alpha}}\subseteq f\left(\gamma\right)\Lambda^{\omega^{\alpha}}$
Thus $g$ is eventually constant. Select $\delta\in\left[0,L\left(f\right)\right)$ such that for $\beta\geq\gamma\geq\delta$, $g\left(\beta\right)=g\left(\gamma\right)$. Since $v$ is $\alpha$ source-regular, $g\left(\delta\right)\not=0$, so choose $e\in f\left(\delta\right)\Lambda^{\omega^{\alpha}}$. We claim $f\leq e$. For $\beta<\delta$, we have $f\left(\beta\right)=f\left(\delta\right)_{\beta}=e_{\beta}$. Suppose $\beta\geq\delta$ and $f\left(\beta\right)\not=e_{\beta}$. Then $e\in f\left(\delta\right)\Lambda^{\omega^{\alpha}}\backslash f\left(\beta\right)\Lambda^{\omega^{\alpha}}$, and $g\left(\beta\right)<g\left(\delta\right)$. This contradicts the construction of $\delta$, so indeed $f\left(\beta\right)=e_{\beta}$. Since $f\leq e$ and $f\in\Lambda^{\infty}$, $e=f$, and $L\left(f\right)=L\left(e\right)=d\left(e\right)+1=\omega^{\alpha}+1>\omega^{\alpha}$. ∎
Theorem 5.8.
If $\Lambda$ is an ordinal graph and $\mathcal{O}\left(\Lambda\right)$ has generators $\left\{T_{e}:e\in\Lambda\right\}$, then there is a representation $\lambda:\mathcal{O}\left(\Lambda\right)\rightarrow B\left(\ell^{2}\left(\Lambda^{\infty}\right)\right)$ such that
$\lambda\left(T_{e}\right)\xi_{f}=\begin{cases}\xi_{ef}&s(e)=r(f)\\ 0&\text{otherwise}\end{cases}$
In particular, $\lambda\left(T_{e}\right)\not=0$ for all $e\in\Lambda$.
Proof.
We will construct a family of operators $\mathcal{V}=\left\{V_{e}:e\in\Lambda\right\}$ satisfying the relations in 4.18. Define $V_{e}\xi_{f}$ by the formula above, i.e. $V_{e}\xi_{f}=\xi_{ef}$ if $s(e)=r(f)$ and $V_{e}\xi_{f}=0$ otherwise. For $f,g\in\Lambda^{\infty}$,
$\left\langle V_{e}\xi_{f},\xi_{g}\right\rangle=\left\langle\xi_{f},V_{e}^{*}\xi_{g}\right\rangle=\begin{cases}1&s(e)=r(f)\text{ and }g=ef\\ 0&\text{otherwise}\end{cases}$
Thus we get the following formula for $V_{e}^{*}$.
$V_{e}^{*}\xi_{g}=\begin{cases}\xi_{f}&g=ef\\ 0&g\not\in e\Lambda^{\infty}\end{cases}$
Now we will verify the relations in 4.18. For relation 1, we have $V_{e}^{*}V_{e}\xi_{f}=0$ if $r(f)\not=s(e)$ and $V_{e}^{*}V_{e}\xi_{f}=\xi_{f}$ if $r(f)=s(e)$, so $V_{e}^{*}V_{e}=V_{s\left(e\right)}$. Relation 2 follows from 5.6. For relations 3 and 4, note that
$V_{e}V_{e}^{*}\xi_{g}=\begin{cases}\xi_{f}&g=eh\text{ for some }h\in\Lambda^{\infty}\\ 0&\text{otherwise}\end{cases}$
If $f\in e\Lambda$ and $g=fh$ for $h\in\Lambda^{\infty}$, then $g=e\left(f^{d(e)}h\right)$ by 5.6. Hence $V_{e}V_{e}^{*}V_{f}V_{f}^{*}=V_{f}V_{f}^{*}$, which implies relation 3. Finally, to verify relation 4 let $v\in\Lambda_{0}$ be an $\alpha$-regular vertex. By 5.7, for every $f\in\Lambda^{\infty}$ with $r\left(f\right)=v$, there exists $g\in v\Lambda^{\omega^{\alpha}}$ such that $f\in g\Lambda^{\infty}$. Hence
$V_{v}=\sum_{g\in v\Lambda^{\omega^{\alpha}}}V_{g}V_{g}^{*}$
Since $\left\{V_{e}:e\in\Lambda\right\}$ satisfies the relations of 4.18, there exists a *-homomorphism $\lambda:\mathcal{O}\left(\Lambda\right)\rightarrow B\left(\ell^{2}\left(\Lambda^{\infty}\right)\right)$ with $\lambda\left(T_{e}\right)=V_{e}$. Then 5.3 shows that for every $e\in\Lambda$ there is $f\in\Lambda^{\infty}$ such that $r\left(f\right)=s\left(e\right)$, and hence $\lambda\left(T_{e}\right)\xi_{f}=\xi_{ef}\not=0$. ∎
Corollary 5.9.
If $\Lambda$ is an ordinal graph and $\mathcal{O}\left(\Lambda\right)$ is generated by $\left\{T_{e}:e\in\Lambda\right\}$, then for every $\alpha\in\mathrm{Ord}$, the map $e\mapsto T_{e}$ is injective on $\Lambda^{\alpha}$.
Proof.
Let $e,f\in\Lambda^{\alpha}$. If $e=fg$, then $d\left(e\right)=d\left(f\right)+d\left(g\right)$. Subtracting $d\left(e\right)$ from both sides, we have $d\left(g\right)=0$, hence $g=s\left(f\right)$ and $e=f$. Thus if $e\not=f$, 3.13 implies $e\Lambda\cap f\Lambda=\emptyset$. By 3.19, we obtain $T_{e}^{*}T_{f}=0$. Then $T_{e}\not=T_{f}$, otherwise $T_{e}^{*}T_{f}=T_{e}^{*}T_{e}=T_{s\left(e\right)}\not=0$. ∎
6. Correspondences
In this section, we show that $\mathcal{O}\left(\Lambda\right)$ may be represented using $\mathrm{C}^{*}$-correspondences in a manner which generalizes the case for directed graphs. Instead of one correspondence, we construct infinitely many correspondences $X_{\alpha}$ over the algebras $\rho_{\alpha}^{\alpha+1}\left(\mathcal{O}\left(\Lambda_{\alpha}\right)\right)$ defined in 3.23. The correspondences are all related by 7.15; in particular, when $\Lambda$ is well-behaved, each $X_{\alpha+1}$ is a correspondence over the Cuntz-Pimsner algebra $\mathcal{O}\left(X_{\alpha}\right)$.
Theorem 6.1.
Let $\alpha\in\mathrm{Ord}$ be fixed and $\mathcal{A}_{\alpha}=\rho_{\alpha}^{\alpha+1}\left(\mathcal{O}\left(\Lambda_{\alpha}\right)\right)\subseteq\mathcal{O}\left(\Lambda_{\alpha+1}\right)$. Suppose $\mathcal{O}\left(\Lambda_{\alpha+1}\right)$ is generated by $\left\{T_{e}:e\in\Lambda_{\alpha}\right\}$. Define
$X_{\alpha}=\overline{\mathrm{span}}\left\{T_{e}T_{f}^{*}\in\mathcal{O}\left(\Lambda_{\alpha+1}\right):\omega^{\alpha}\leq d\left(e\right)<\omega^{\alpha}\cdot 2,f\in\Lambda_{\alpha}\right\}$
For $x,y\in X_{\alpha}$ and $a\in\mathcal{A}_{\alpha}$, define
$x\cdot a$ $=xa\in X_{\alpha}$
$\varphi_{\alpha}\left(a\right)x$ $=a\cdot x=ax\in X_{\alpha}$
$\left\langle x,y\right\rangle_{\alpha}$ $=x^{*}y\in\mathcal{A}_{\alpha}$
where $\varphi_{\alpha}:\mathcal{A}_{\alpha}\rightarrow\mathcal{L}\left(X_{\alpha}\right)$. Then $\left(X_{\alpha},\varphi_{\alpha}\right)$ is a correspondence over $\mathcal{A}_{\alpha}$.
Proof.
Fix $\alpha\in\mathrm{Ord}$. Define
[…]
By 3.21, ¯$\overline{\mathcal{B}}=\mathcal{A}_{\alpha}$, and we also have $\overline{Y}=X_{\alpha}$. First we show that $Y$ is closed under the right action of $\mathcal{B}$. If $x=T_{e}T_{f}^{*}$ for some $\omega^{\alpha}\leq d\left(e\right)<\omega^{\alpha}\cdot 2$ and $f\in\Lambda_{\alpha}$, choose $g,h\in\Lambda$ such that $d\left(g\right)=\omega^{\alpha}$ and $e=gh$. Then $x=T_{g}T_{h}T_{f}^{*}$ for $h,f\in\Lambda_{\alpha}$. If $a=T_{p}T_{q}^{*}$ for $p,q\in\Lambda_{\alpha}$, then by 3.19, $x\cdot a=T_{g}T_{h}T_{f}^{*}T_{p}T_{q}^{*}$ is either 00 or $T_{g}T_{r}T_{s}^{*}$ for some $r,s\in\Lambda_{\alpha}$. Since $\mathcal{B}$ and $Y$ are the span of elements of these forms, it follows that $Y$ is closed under the right action of $\mathcal{B}$.
Next we will prove the function $\left(x,y\right)\mapsto x^{*}y=\left\langle x,y\right\rangle_{\alpha}$ with domain $Y\times Y$ maps into $\mathcal{B}$. Let $x=T_{e}T_{f}^{*}$ and $y\in T_{g}T_{h}^{*}$ with $\omega^{\alpha}\leq\min\left\{d\left(e\right),d\left(g\right)\right\}$, $\max\left\{d\left(e\right),d\left(g\right)\right\}<\omega^{\alpha}\cdot 2$, and $f,h\in\Lambda_{\alpha}$. Then
$\left\langle x,y\right\rangle_{\alpha}=\left(T_{e}T_{f}^{*}\right)^{*}T_{g}T_{h}^{*}=T_{f}T_{e}^{*}T_{g}T_{h}^{*}$
As $T_{f},T_{h}^{*}\in\mathcal{B}$ and $\mathcal{B}$ is a linear subspace, it suffices to show $T_{e}^{*}T_{g}\in\mathcal{B}$. By taking adjoints, we may assume without loss of generality $d\left(e\right)\leq d\left(g\right)$. Then 3.19 implies that either $T_{e}^{*}T_{g}=0$, or there exists $p\in\Lambda$ such that $g=ep$ and $T_{e}^{*}T_{g}=T_{p}$. If $g=ep$, then $\omega^{\alpha}\leq d\left(g\right)=d\left(e\right)+d\left(p\right)<\omega^{\alpha}\cdot 2$. Subtracting $\omega^{\alpha}$ from both sides, we see $-\omega^{\alpha}+d\left(e\right)+d\left(p\right)<-\omega^{\alpha}+\omega^{\alpha}\cdot 2=\omega^{\alpha}$, and in particular, $d\left(p\right)<\omega^{\alpha}$. Thus $T_{e}^{*}T_{g}=T_{p}\in\mathcal{B}$, as desired.
The operation $\left\langle\cdot,\cdot\right\rangle_{\alpha}$ is easily verified to be an inner product on $Y$. Moreover, the inner product on $Y$ induces a norm which agrees with the C*-norm, as the following calculation shows.
$\left\|x\right\|_{Y}=\left\|\left\langle x,x\right\rangle\right\|_{\mathcal{O}\left(\Lambda_{\alpha+1}\right)}^{1/2}=\left\|x^{*}x\right\|_{\mathcal{O}\left(\Lambda_{\alpha+1}\right)}^{1/2}=\left\|x\right\|_{\mathcal{O}\left(\Lambda_{\alpha+1}\right)}$
Additionally, the right action of $\mathcal{B}$ is continuous, hence extends to a right action of $\mathcal{A}_{\alpha}$ on $X_{\alpha}$ which agrees with right multiplication in $\mathcal{O}\left(\Lambda_{\alpha+1}\right)$. Similarly, the $\mathcal{B}$-valued inner product on $Y$ extends to an inner product on $X_{\alpha}$ valued in $\mathcal{A}_{\alpha}$.
All that is left is to verify that $\varphi_{\alpha}$ is a *-homomorphism into $\mathcal{L}\left(X_{\alpha}\right)$. The following calculation does this.
$\left\langle\varphi_{\alpha}\left(a\right)x,y\right\rangle_{\alpha}$ $=\left(\varphi_{\alpha}\left(a\right)x\right)^{*}y=\left(ax\right)^{*}y$
$=x^{*}a^{*}y=x^{*}\left(\varphi_{\alpha}\left(a^{*}\right)y\right)=\left\langle x,\varphi_{\alpha}\left(a^{*}\right)y\right\rangle_{\alpha}$
Example 6.2.
Consider $X_{\alpha}$ for $\alpha=0$. Then $\mathcal{A}_{0}=\rho_{0}^{1}\left(\mathcal{O}\left(\Lambda_{0}\right)\right)$, and $\mathcal{O}\left(\Lambda_{0}\right)=c_{0}\left(\Lambda_{0}\right)$. Since for each $v\in\Lambda_{0}$ we have $T_{v}\not=0$, $\rho_{0}^{1}$ is injective. In particular, $\mathcal{A}_{0}$ is isomorphic to $c_{0}\left(\Lambda_{0}\right)$. Moreover, if $e,f\in\Lambda$ and $d\left(f\right)=0$, $T_{e}T_{f}^{*}=T_{e}T_{f}\in\left\{0,T_{e}\right\}$. Thus $X_{0}=\overline{\mathrm{span}}\left\{T_{e}:e\in\Lambda^{1}\right\}$, and by 5.9 the elements of $\left\{T_{e}:e\in\Lambda^{1}\right\}$ are distinct. In the above proof, we have $Y=\mathrm{span}\left\{T_{e}:e\in\Lambda^{1}\right\}$ is isomorphic as an inner product $c_{0}\left(\Lambda_{0}\right)$-module to $c_{c}\left(\Lambda^{1}\right)$, where for $a\in c_{0}\left(\Lambda_{0}\right)$ and $x,y\in c_{c}\left(\Lambda^{1}\right)$,
$\left(x\cdot a\right)\left(e\right)$ $=x\left(e\right)a\left(s\left(e\right)\right)\text{ for }e\in\Lambda^{1}$
$\left\langle x,y\right\rangle\left(v\right)$ $=\sum_{e\in\Lambda^{1}v}\overline{x\left(e\right)}y\left(e\right)\text{ for }v\in\Lambda^{0}$
Moreover, there is a left action $\phi_{\alpha}:c_{0}\left(\Lambda_{0}\right)\rightarrow\mathcal{L}\left(c_{c}\left(\Lambda^{1}\right)\right)$ defined by
$\left(\phi_{\alpha}\left(a\right)x\right)\left(e\right)=\left(a\cdot x\right)\left(e\right)=a\left(r\left(e\right)\right)x\left(e\right)\text{ for }e\in\Lambda^{1}$
which agrees with $\varphi_{\alpha}$ on $Y$. Thus by [6, Proposition 3.10], $\mathcal{O}\left(X_{0}\right)\cong\mathcal{O}\left(\Lambda_{1}\right)$.
Definition 6.3.
Let $\pi_{\alpha}:\mathcal{A}_{\alpha}\rightarrow\mathcal{O}\left(\Lambda_{\alpha+1}\right)$ be the inclusion, which is a *-homomorphism. Define $\psi_{\alpha}:X_{\alpha}\rightarrow\mathcal{O}\left(\Lambda_{\alpha}\right)$ to be the inclusion, which is an isometry.
By the definition of $X_{\alpha}$, $\left(\psi_{\alpha},\pi_{\alpha}\right)$ is a Toeplitz representation of $X_{\alpha}$. Understanding the Katsura ideal will be key to proving 7.15. We recall the definition below.
Definition 6.4 ([7, Definition 2.3]).
The Katsura ideal $J_{\alpha}$ is defined as
$J_{\alpha}=J_{X_{\alpha}}=\varphi_{\alpha}^{-1}\left(\mathcal{K}\left(X_{\alpha}\right)\right)\cap\left(\ker\varphi_{\alpha}\right)^{\perp}$
When $\Lambda$ is a directed graph, the Katsura ideal $J_{0}=c_{0}\left(\left\{v\in\Lambda_{0}:0<\left|r^{-1}\left(v\right)\right|<\infty\right\}\right)$ is the ideal generated by the vertex projections of 00 regular vertices. We would like know in general whether $J_{\alpha}$ is generated by the vertex projections $\left\{T_{v}:v\text{ is }\alpha\text{ regular}\right\}$. Unfortunately, this is more difficult to determine for $\alpha>0$ since the algebras $\mathcal{A}_{\alpha}$ are not necessarily commutative. The following result, which will be important later, makes some progress towards answering this question.
Proposition 6.5.
$v\in\Lambda_{0}$ is $\alpha$ regular if and only if $T_{v}\in J_{\alpha}$.
Proof.
Let $\left\{T_{e}:e\in\Lambda_{\alpha+1}\right\}$ be the generators of $\mathcal{O}\left(\Lambda_{\alpha+1}\right)$. Then $\mathcal{A}_{\alpha}$ is generated by $\left\{T_{e}:e\in\Lambda_{\alpha}\right\}$. Suppose $T_{v}\in J_{\alpha}$ and $e\in\Lambda_{\alpha}$ with $r(e)=v$ such that $s(e)$ is an $\alpha$ source. Then for $f\in\Lambda^{\omega^{\alpha}}$, $T_{e}T_{f}=0$. Since $Y_{\alpha}$ is the $\mathcal{B}_{\alpha}$-span of such $T_{f}$, $T_{e}\in\ker\varphi_{\alpha}$. However, $T_{v}T_{e}=T_{e}$, which is non-zero by 5.8. This contradicts $T_{v}\in\left(\ker\varphi_{\alpha}\right)^{\perp}$, so $v$ is $\alpha$ source-regular.
To see that $v$ is $\alpha$ row-finite, suppose $v\Lambda^{\omega^{\alpha}}$ is infinite. Then for each $f\in v\Lambda^{\omega^{\alpha}}$, $\varphi_{\alpha}\left(T_{v}\right)T_{f}=T_{f}$. We will show that this contradicts the compactness of $\varphi_{\alpha}\left(T_{v}\right)$. Since $\varphi_{\alpha}\left(T_{v}\right)$ is compact, choose $m\in\mathbb{N}$ and $x_{n},y_{n}\in X_{\alpha}$ such that
$\left\|\varphi_{\alpha}\left(T_{v}\right)-\sum_{n=1}^{m}\theta_{x_{n},y_{n}}\right\|<\frac{1}{2}$
For fixed $x\in X_{\alpha}$, the map $y\mapsto\theta_{x,y}$ is continuous in the operator norm topology, as shown by the following estimate:
$\left\|\left(\theta_{x,y}-\theta_{x,y^{\prime}}\right)z\right\|=\left\|x\cdot\left\langle y-y^{\prime},z\right\rangle\right\|\leq\left\|x\right\|\left\|z\right\|\left\|y-y^{\prime}\right\|$
[…]
$\left\|\sum_{n=1}^{m}\left(\theta_{x_{n},y_{n}}-\theta_{x_{n},y_{n}^{\prime}}\right)\right\|<\sum_{n=1}^{m}\frac{\left\|x_{n}\right\|}{2m\left\|x_{n}\right\|}=\frac{1}{2}$
For $g\in\Lambda^{\omega^{\alpha}}$, $\left\langle T_{e}T_{f}^{*},T_{g}\right\rangle=T_{f}T_{e}^{*}T_{g}$ is non-zero only if $e_{\omega^{\alpha}}=g$. Each $y_{n}^{\prime}$ is a finite linear combination of elements of the form $T_{e}T_{f}^{*}$, and in particular, there are only finitely many $g\in\Lambda^{\omega^{\alpha}}$ with $\left\langle y_{n}^{\prime},T_{g}\right\rangle\not=0$. Choose $g\in v\Lambda^{\omega^{\alpha}}$ such that $\theta_{x_{n},y_{n}^{\prime}}T_{g}=0$ for all $1\leq n\leq m$. Then
$\left\|\varphi_{\alpha}\left(T_{v}\right)T_{g}-\sum_{n=1}^{m}\theta_{x_{n},y_{n}}T_{g}\right\|$ $=\left\|T_{g}-\sum_{n=1}^{m}\left(\theta_{x_{n},y_{n}}-\theta_{x_{n},y_{n}^{\prime}}\right)T_{g}\right\|$
$\geq\left\|T_{g}\right\|-\left\|\sum_{n=1}^{m}\left(\theta_{x_{n},y_{n}}-\theta_{x_{n},y_{n}^{\prime}}\right)T_{g}\right\|>1-\frac{1}{2}=\frac{1}{2}$
This is a contradiction, hence $v$ is $\alpha$ row-finite.
Suppose for the converse $v$ is $\alpha$ regular. Then $T_{v}=\sum_{f\in v\Lambda^{\omega^{\alpha}}}T_{f}T_{f}^{*}$. If $a\in\ker\varphi_{\alpha}$, this implies
$aT_{v}=a\sum_{f\in v\Lambda^{\omega^{\alpha}}}T_{f}T_{f}^{*}=\sum_{f\in v\Lambda^{\omega^{\alpha}}}aT_{f}T_{f}^{*}=0$
Therefore $T_{v}\in\left(\ker\varphi_{\alpha}\right)^{\perp}$. Moreover, for $x\in X_{\alpha}$,
$\varphi_{\alpha}\left(T_{v}\right)x=T_{v}x=\sum_{f\in v\Lambda^{\omega^{\alpha}}}T_{f}T_{f}^{*}x=\sum_{f\in v\Lambda^{\omega^{\alpha}}}T_{f}\cdot\left\langle T_{f},x\right\rangle=\sum_{f\in v\Lambda^{\omega^{\alpha}}}\theta_{T_{f},T_{f}}x$
From this we see $\varphi_{\alpha}\left(T_{v}\right)=\sum_{f\in v\Lambda^{\omega^{\alpha}}}\theta_{T_{f},T_{f}}$, and $T_{v}\in\left(\ker\varphi_{\alpha}\right)^{\perp}\cap\varphi_{\alpha}^{-1}\left(\mathcal{K}\left(X_{\alpha}\right)\right)=J_{\alpha}$. ∎
7. Cuntz-Krieger Uniqueness
Our next goal is to identify conditions under which a *-homomorphism $\pi:\mathcal{O}\left(\Lambda\right)\rightarrow\mathcal{C}$ is injective. Our strategy is to induct on $\alpha$ and prove $\pi$ restricted to the algebras generated by $\left\{T_{e}:e\in\Lambda_{\alpha}\right\}$ is injective. At each step of the induction we will apply [4, Theorem 3.9]. First we review the definition of a non-returning vector for a correspondence.
Definition 7.1 ([4, Definition 3.1]).
If $X$ is a $\mathrm{C}^{*}$-correspondence over $A$ and $\left(\psi_{X},\pi_{A}\right):\left(X,A\right)\rightarrow\mathcal{O}\left(X\right)$ is the universal covariant representation, then $\zeta\in X^{\otimes m}$ is non-returning if for all $1\leq n<m$ and $\xi\in X^{\otimes n}$,
$\psi_{X}^{\otimes m}\left(\zeta\right)^{*}\psi_{X}^{\otimes n}\left(\xi\right)\psi_{X}^{\otimes m}\left(\zeta\right)=0$
Definition 7.2 (cf. [9, Lemma 3.7]).
For $\alpha\in\mathrm{Ord}$ and $n<\omega$, $e\in\Lambda^{\omega^{\alpha}\cdot n}$ is non-returning if for all $f\in\Lambda$ with $\omega^{\alpha}\leq d(f)<\omega^{\alpha}\cdot n$ and $\beta<\omega^{\alpha}$, $fe^{\beta}\not\in e\Lambda$.
In 7.9 we will show every non-returning path of $\Lambda^{\omega^{\alpha}\cdot m}$ is associated with a non-returning vector in $X_{\alpha}^{\otimes m}$.
Definition 7.3 ([4, Condition (S)]).
A $\mathrm{C}^{*}$-correspondence $X$ over $A$ satisfies condition (S) if for all $a\in A$ with $a\geq 0$, all $n\in\mathbb{N}$, and all $\varepsilon>0$ there is $m>n$ and non-returning $\zeta\in X^{\otimes m}$ with $\|\zeta\|=1$ such that
$\left\|\left\langle\zeta,a\zeta\right\rangle\right\|>\left\|a\right\|-\varepsilon$
Returning to the context of ordinal graphs, for $\alpha\in\mathrm{Ord}$, let $\left(\mu_{\alpha},\eta_{\alpha}\right):\left(X_{\alpha},\mathcal{A}_{\alpha}\right)\rightarrow\mathcal{O}\left(X_{\alpha}\right)$ be the universal covariant representation. For the rest of the section, let $\left\{T_{e}:e\in\Lambda_{\alpha+1}\right\}$ be the generators for $\mathcal{O}\left(\Lambda_{\alpha+1}\right)$ and $\left\{S_{e}:e\in\Lambda_{\alpha}\right\}$ be the generators for $\mathcal{O}\left(\Lambda_{\alpha}\right)$.
Definition 7.4.
For $e\in\Lambda$ with $\omega^{\alpha}\cdot n\leq d\left(e\right)<\omega^{\alpha}\cdot(n+1)$, define $\chi_{e}\in X_{\alpha}^{\otimes n}$ so that
$\chi_{e}=T_{f_{1}}\otimes T_{f_{2}}\otimes\ldots\otimes T_{f_{n}}T_{g}$
where $d\left(f_{k}\right)=\omega^{\alpha}$, $d\left(g\right)<\omega^{\alpha}$, and $e=f_{1}f_{2}\ldots f_{n}g$.
By 4.19, the representation of $e$ as $f_{1}f_{2}\dots f_{n}g$ exists and is unique, hence $\chi_{e}$ is well-defined.
Proposition 7.5.
The family $\left\{\mu_{\alpha}^{\otimes n}\left(\chi_{e}\right):n\in\mathbb{N},e\in\Lambda,\omega^{\alpha}\cdot n\leq d(e)<\omega^{\alpha}\cdot(n+1)\right\}$ is a Cuntz-Krieger $\Lambda_{\alpha+1}$-family.
Proof.
We will show that there are operators satisfiying the relations of 4.20. For $e\in\Lambda_{0}\cup\Lambda^{\omega^{\beta}}$ with $\beta<\alpha$, let $U_{e}=\eta_{\alpha}\left(T_{e}\right)$, and if $e\in\Lambda^{\omega^{\alpha}}$, define $U_{e}=\mu_{\alpha}\left(\chi_{e}\right)$. Relations 1 through 4 hold for the generators $\left\{U_{e}:e\in\Lambda_{0}\cup\Lambda^{\omega^{\beta}},\beta<\alpha\right\}$ since $\eta_{\alpha}$ is a *-homomorphism. Thus it sufficies to check the relations for $U_{e}$ with $e\in\Lambda^{\omega^{\alpha}}$. Relation 1 follows because
$U_{e}^{*}U_{e}=\mu_{\alpha}\left(\chi_{e}\right)^{*}\mu_{\alpha}\left(\chi_{e}\right)=\eta_{\alpha}\left(\left\langle\chi_{e},\chi_{e}\right\rangle\right)=\eta_{\alpha}\left(T_{e}^{*}T_{e}\right)=\eta_{\alpha}\left(T_{s(e)}\right)=U_{s\left(e\right)}$
For relation 2, suppose $f\in\Lambda_{0}\cup\Lambda^{\omega^{\beta}}$ for $\beta<\alpha$. Then if $s(f)=r(e)$,
$U_{f}U_{e}=\eta_{\alpha}\left(T_{f}\right)\mu_{\alpha}\left(\chi_{e}\right)=\mu_{\alpha}\left(\varphi_{\alpha}\left(T_{f}\right)\chi_{e}\right)=\mu_{\alpha}\left(\chi_{fe}\right)=U_{fe}$
There are 2 cases for relation 3. First we suppose that $f\in\Lambda_{0}\cup\Lambda^{\omega^{\beta}}$, $\beta<\alpha$, and $e\Lambda\cap f\Lambda=\emptyset$, in which case
$U_{f}^{*}U_{e}=\eta_{\alpha}\left(T_{f}^{*}\right)\mu_{\alpha}\left(\chi_{e}\right)=\mu_{\alpha}\left(\varphi_{\alpha}\left(T_{e}^{*}\right)\chi_{e}\right)=\mu_{\alpha}\left(0\right)=0$
For the other case, suppose $f\in\Lambda^{\omega^{\alpha}}$ and $e\Lambda\cap f\Lambda=\emptyset$. Then
$U_{f}^{*}U_{e}=\mu_{\alpha}\left(\chi_{f}\right)^{*}\mu_{\alpha}\left(\chi_{e}\right)=\eta_{\alpha}\left(\left\langle\chi_{f},\chi_{e}\right\rangle\right)=\eta_{\alpha}\left(T_{f}^{*}T_{e}\right)=\eta_{\alpha}\left(0\right)=0$
Finally, relation 4 follows from covariance of $\left(\mu_{\alpha},\eta_{\alpha}\right)$. In particular by 6.5, if $v$ is $\alpha$ regular, $T_{v}\in J_{\alpha}$. Then by covariance,
$\eta_{\alpha}\left(T_{v}\right)$ $=U_{v}=\left(\mu_{\alpha},\eta_{\alpha}\right)^{(1)}\left(\varphi_{\alpha}\left(T_{v}\right)\right)=\left(\mu_{\alpha},\eta_{\alpha}\right)^{(1)}\left(\sum_{f\in v\Lambda^{\omega^{\alpha}}}\theta_{\chi_{f},\chi_{f}}\right)$
$=\sum_{f\in v\Lambda^{\omega^{\alpha}}}\left(\mu_{\alpha},\eta_{\alpha}\right)^{(1)}\left(\theta_{\chi_{f},\chi_{f}}\right)=\sum_{f\in v\Lambda^{\omega^{\alpha}}}\mu_{\alpha}\left(\chi_{f}\right)\mu_{\alpha}\left(\chi_{f}\right)^{*}=\sum_{f\in v\Lambda^{\omega^{\alpha}}}U_{f}U_{f}^{*}$
Definition 7.6.
Let $j_{\alpha}:\mathcal{O}\left(\Lambda_{\alpha+1}\right)\rightarrow\mathcal{O}\left(X_{\alpha}\right)$ be the surjective *-homomorphism induced by the universal property of $\mathcal{O}\left(\Lambda_{\alpha+1}\right)$.
Note then that for $e\in\Lambda^{\omega^{\alpha}}$, $j_{\alpha}\left(T_{e}\right)=\mu_{\alpha}\left(\chi_{e}\right)$, and for $e\in\Lambda_{\alpha}$, $j_{\alpha}\left(T_{e}\right)=\eta_{\alpha}\left(T_{e}\right)$.
Corollary 7.7.
If $n,m\in\mathbb{N}$ with $\omega^{\alpha}\cdot n\leq d(e)<\omega^{\alpha}\cdot(n+1)$ and $\omega^{\alpha}\cdot m\leq d(f)<\omega^{\alpha}\cdot(m+1)$, then
$\mu_{Y_{\alpha}}^{\otimes n}\left(\chi_{e}\right)^{*}\mu_{Y_{\alpha}}^{\otimes m}\left(\chi_{f}\right)=\begin{cases}\mu_{\alpha}^{\otimes(m-n)}\left(\chi_{g}\right)&f=eg\\ \mu_{\alpha}^{\otimes(n-m)}\left(\chi_{g}\right)^{*}&e=fg\\ 0&e\Lambda\cap f\Lambda=\emptyset\end{cases}$
Proof.
Fix $n,m\in\mathbb{N}$ with $\omega^{\alpha}\cdot n\leq d(e)<\omega^{\alpha}\cdot(n+1)$ and $\omega^{\alpha}\cdot m\leq d(f)<\omega^{\alpha}\cdot(m+1)$. By 3.19,
$j_{\alpha}\left(T_{e}^{*}T_{f}\right)$ $=j_{\alpha}\left(T_{e}\right)^{*}j_{\alpha}\left(T_{f}\right)=\mu_{\alpha}^{\otimes n}\left(\chi_{e}\right)^{*}\mu_{\alpha}^{\otimes m}\left(\chi_{f}\right)$
=$=\begin{cases}j_{\alpha}\left(T_{g}\right)&f=eg\\ j_{\alpha}\left(T_{g}\right)^{*}&e=fg\\ 0&e\Lambda\cap f\Lambda=\emptyset\end{cases}$
=$=\begin{cases}\mu_{\alpha}^{\otimes(m-n)}\left(\chi_{g}\right)&f=eg\\ \mu_{\alpha}^{\otimes(n-m)}\left(\chi_{g}\right)^{*}&e=fg\\ 0&e\Lambda\cap f\Lambda=\emptyset\end{cases}$
Lemma 7.8.
Let $(X,\varphi)$ be a $C^{*}$-correspondence over $A$ with universal covariant representation $\left(\psi_{X},\pi_{A}\right):\left(X,A\right)\rightarrow\mathcal{O}\left(X\right)$. If $\zeta\in X^{\otimes m}$ is non-returning, then for all $1\leq n<m$, $\xi\in X^{\otimes n}$, and $a,a^{\prime},b,b^{\prime}\in A$,
$\psi_{X}^{\otimes m}\left(\varphi_{m}\left(a\right)\zeta b\right)^{*}\psi_{X}^{\otimes n}\left(\xi\right)\psi_{X}^{\otimes m}\left(\varphi_{m}\left(a^{\prime}\right)\zeta b^{\prime}\right)=0$
Proof.
Let $1\leq n<m$ and $\xi\in X^{\otimes n}$ be given. Then since $\varphi_{n}\left(a^{*}\right)\xi a^{\prime}\in X^{\otimes n}$,
$\psi_{X}^{\otimes m}\left(\varphi_{m}(a)\zeta b\right)^{*}\psi_{X}^{\otimes n}\left(\xi\right)\psi_{X}^{\otimes m}\left(\varphi_{m}(a^{\prime})\zeta b^{\prime}\right)$
$=$ $\pi_{A}\left(b\right)^{*}\psi_{X}^{\otimes m}\left(\zeta\right)^{*}\pi_{A}\left(a\right)^{*}\psi_{X}^{\otimes n}\left(\xi\right)\pi_{A}\left(a^{\prime}\right)\psi_{X}^{\otimes m}\left(\zeta\right)\pi_{A}\left(b^{\prime}\right)$
$=$ $\pi_{A}\left(b\right)^{*}\psi_{X}^{\otimes m}\left(\zeta\right)^{*}\psi_{X}^{\otimes n}\left(\varphi_{n}\left(a^{*}\right)\xi a^{\prime}\right)\psi_{X}^{\otimes m}\left(\zeta\right)\pi_{A}\left(b^{\prime}\right)$
$=$ $\pi_{A}\left(b\right)^{*}0\pi_{A}\left(b^{\prime}\right)=0$
Lemma 7.9.
If $e\in\Lambda^{\omega^{\alpha}\cdot m}$ is non-returning, then $\chi_{e}\in X_{\alpha}^{\otimes m}$ is non-returning.
Proof.
Let $1\leq n<m$ be given. Since $X_{\alpha}^{\otimes n}$ is generated by
$\left\{\chi_{f}\cdot T_{g}^{*}:g\in\Lambda_{\alpha},f\in\Lambda,\omega^{\alpha}\cdot n\leq d(f)<\omega^{\alpha}\cdot(n+1)\right\}$
it suffices to check $\mu_{\alpha}^{\otimes m}\left(\chi_{e}\right)^{*}\mu_{\alpha}^{\otimes n}\left(\chi_{f}\cdot T_{g}^{*}\right)\mu_{\alpha}^{\otimes m}\left(\chi_{e}\right)=0$. If
$\mu_{\alpha}^{\otimes n}\left(\chi_{f}\cdot T_{g}^{*}\right)\mu_{\alpha}^{\otimes m}\left(\chi_{e}\right)=\mu_{\alpha}^{\otimes n}\left(\chi_{f}\right)\mu_{\alpha}^{\otimes m}\left(\varphi_{\alpha}\left(T_{g}^{*}\right)\chi_{e}\right)\not=0$
then $e\in g\Lambda$, $s(f)=r\left(e^{d(g)}\right)$, and
$\mu_{\alpha}^{\otimes n}\left(\chi_{f}\right)\mu_{\alpha}^{\otimes m}\left(\chi_{e^{d(g)}}\right)=\mu_{\alpha}^{\otimes\left(n+m\right)}\left(\chi_{fe^{d(g)}}\right)$
Since $e$ is non-returning, $fe^{d(g)}\not\in e\Lambda$. Thus by 7.7,
$\mu_{\alpha}^{\otimes m}\left(\chi_{e}\right)^{*}\mu_{\alpha}^{\otimes n}\left(\chi_{f}\cdot T_{g}^{*}\right)\mu_{\alpha}^{\otimes m}\left(\chi_{e}\right)=\mu_{\alpha}^{\otimes m}\left(\chi_{e}\right)^{*}\mu_{\alpha}^{\otimes(n+m)}\left(\chi_{fe^{d(g)}}\right)=0$
Lemma 7.10.
Let $\left(X,\varphi\right)$ be a correspondence over a $C^{*}$-algebra $A$. If $\zeta\in X^{\otimes m}$ is non-returning, then
$K=\overline{\text{$\mathrm{span}$}}\left\{\varphi\left(a\right)\zeta\cdot b:a,b\in A\right\}$
is a closed Hilbert submodule of $X^{\otimes m}$, and each $\eta\in K$ is non-returning.
Proof.
Let $\left(\psi,\pi\right):\left(X,\varphi\right)\rightarrow\mathcal{O}\left(X\right)$ be the universal covariant representation of $X$. Define $Y=\text{$\mathrm{span}$}\left\{\varphi\left(a\right)\zeta\cdot b:a,b\in A\right\}$. Then $Y$ is closed under the right action of $A$, hence $Y$ is a submodule of $X$. Since the right action is continuous, $\overline{Y}=K$ is a closed Hilbert submodule of $X$. For $1\leq n<m$ and $\xi\in X^{\otimes n}$, the function
$\nu\mapsto\psi^{\otimes m}\left(\nu\right)^{*}\psi^{\otimes n}\left(\xi\right)\psi^{\otimes m}\left(\nu\right)$
is continuous. Thus the set of non-returning vectors in $X^{\otimes m}$ is closed, and it suffices to show that every vector in $Y$ is non-returning. Let $\kappa\in Y$ be arbitrary, and choose $a_{k},b_{k}\in A$ with
$\kappa=\sum_{k=1}^{r}\varphi\left(a_{k}\right)\zeta\cdot b_{k}$
Since we wish to show $\kappa$ is non-returning, let $1\leq n<m$ and $\xi\in X^{\otimes n}$ be given. Then
$\psi^{\otimes m}\left(\kappa\right)^{*}\psi^{\otimes n}\left(\xi\right)\psi^{\otimes m}\left(\kappa\right)=\sum_{j,k=1}^{r}\psi^{\otimes m}\left(\varphi\left(a_{j}\right)\zeta\cdot b_{j}\right)^{*}\psi^{\otimes n}\left(\xi\right)\psi^{\otimes m}\left(\varphi\left(a_{k}\right)\zeta\cdot b_{k}\right)$
By 7.8, this is 00, and $\kappa$ is non-returning. ∎
We are now in the position to prove a version of the Cuntz-Krieger uniqueness theorem for ordinal graphs. To do so, we introduce what we call condition (S) for ordinal graphs, which represents the circumstance in which each correspondence $X_{\alpha}$ satisfies condition (S).
Definition 7.11.
A path $e\in\Lambda$ is $\alpha$-full if $e\in\Lambda\backslash\Lambda_{\alpha}$ and for every $v\in\Lambda_{0}$ in the same connected component of $\Lambda_{\alpha}$ as $r\left(e\right)$, there exists $\beta<\omega^{\alpha}$ and $f\in\Lambda_{\alpha}$ such that $s\left(f\right)=r\left(e^{\beta}\right)$ and $r\left(f\right)=v$.
Remark 7.12.
Since the connected componenets of $\Lambda_{0}$ are the vertices, every path $e\in\Lambda\backslash\Lambda_{0}$ is 00-full. In particular, all edges in a directed graph are 00-full.
Definition 7.13.
An ordinal graph $\Lambda$ satisfies condition (S) if for every $\alpha\in\mathrm{Ord}$ such that $\Lambda^{\omega^{\alpha}}\not=\emptyset$, every connected component $F$ of $\Lambda_{\alpha}$, and every $n\in\mathbb{N}$ there exists non-returning, $\alpha$-full $f\in\Lambda$ with $r\left(f\right)\in F$ and $\omega^{\alpha}\cdot n\leq d\left(f\right)<\omega^{\alpha+1}$.
Remark 7.14.
If $\Lambda$ satisfies condition (S) and $\Lambda^{\omega^{\alpha}}\not=\emptyset$, then for each $v\in\Lambda_{0}$ there exists $f\in\Lambda_{\alpha}$, $\alpha$-full $e\in\Lambda$, and $\beta<\omega^{\alpha}$ such that $s\left(f\right)=r\left(e^{\beta}\right)$ and $r\left(f\right)=v$. Then $r\left(fe^{\beta}\right)=v$ and $d\left(fe^{\beta}\right)=d\left(f\right)-\beta+d\left(e\right)\geq\omega^{\alpha}$. Thus $\Lambda$ has no $\alpha$-sources.
Theorem 7.15 (Cuntz-Krieger Uniqueness Theorem for Ordinal Graphs).
If $\Lambda$ is an ordinal graph with no 1-regular vertices which satisfies condition (S) and $\Lambda^{\omega^{\alpha}}\not=\emptyset$, then
  1. (1)
    $J_{\alpha}=0$ if $\alpha\geq 1$.
  2. (2)
    $\mathcal{O}\left(X_{\alpha}\right)\cong\mathcal{O}\left(\Lambda_{\alpha+1}\right)$.
  3. (3)
    $X_{\alpha}$ satisfies condition (S).
  4. (4)
    If $\pi:\mathcal{O}\left(\Lambda_{\alpha+1}\right)\rightarrow\mathcal{C}$ is a *-homomorphism into a $C^{*}$-algebra $\mathcal{C}$ such that for every $v\in\Lambda_{0}$, $\pi\left(T_{v}\right)\not=0$, then $\pi$ is injective.
In particular, if $\Lambda$ satisfies condition (S) and has no 1-regular vertices, then a *-homomorphism $\pi:\mathcal{O}\left(\Lambda\right)\rightarrow\mathcal{C}$ is injective iff $\pi\left(T_{v}\right)\not=0$ for all $v\in\Lambda_{0}$.
Proof.
The proof is by transfinite induction. Let $\alpha$ with $\Lambda^{\omega^{\alpha}}$ be fixed, $\left\{S_{e}:e\in\Lambda_{\alpha}\right\}$ be the generators of $\mathcal{O}\left(\Lambda_{\alpha}\right)$, and $\left\{T_{e}:e\in\Lambda_{\alpha+1}\right\}$ be the generators of $\mathcal{O}\left(\Lambda_{\alpha+1}\right)$. Suppose we have Cuntz-Krieger uniqueness for $\alpha$, or more specifically, that for every ordinal graph $\Gamma$ with $\mathcal{O}\left(\Gamma_{\alpha}\right)$ generated by $\left\{W_{e}:e\in\Gamma_{\alpha}\right\}$ and every *-homomorphism $\pi:\mathcal{O}\left(\Gamma_{\alpha}\right)\rightarrow\mathcal{C}$ with $\pi\left(W_{v}\right)\not=0$ for all $v\in\Gamma_{0}$, $\pi$ is injective. In particular, we have Cuntz-Krieger uniqueness for $\Lambda_{\alpha}$. Note that when $\alpha=0$, $\mathcal{O}\left(\Lambda_{\alpha}\right)\cong c_{0}\left(\Lambda_{0}\right)$ has Cuntz-Krieger uniqueness, and this is the base case. First we will prove statements 1, 2, and 3 in the theorem, and then prove that $\mathcal{O}\left(\Lambda_{\alpha+1}\right)$ has the Cuntz-Krieger uniqueness property. Afterwards, we will prove that $\mathcal{O}\left(\Lambda_{\beta}\right)$ has Cuntz-Krieger uniqueness when $\beta$ is a limit ordinal, which will complete the proof.
To see 1, suppose $\alpha\geq 1$. Note that $\Lambda$ has no $1$-regular vertices, so by 6.5 and 4.6, $T_{v}\not\in J_{\alpha}$ for every $v\in\Lambda_{0}$. Let $\mathcal{I}=\left(\rho_{\alpha}^{\alpha+1}\right)^{-1}\left(J_{\alpha}\right)\vartriangleleft\mathcal{O}\left(\Lambda_{\alpha}\right)$, and $q:\mathcal{O}\left(\Lambda_{\alpha}\right)\rightarrow\mathcal{O}\left(\Lambda_{\alpha}\right)/\mathcal{I}$ be the quotient map. Since $\rho_{\alpha}^{\alpha+1}\left(S_{v}\right)=T_{v}$, $S_{v}\not\in\mathcal{I}$ for all $v\in\Lambda_{0}$. Then $q\left(S_{v}\right)\not=0$ for all $v\in\Lambda_{0}$, and by hypothesis, $q$ is injective. Therefore $\mathcal{I}=0$, and since $J_{\alpha}\subseteq\rho_{\alpha}^{\alpha+1}\left(\mathcal{O}\left(\Lambda_{\alpha}\right)\right)$, $J_{\alpha}=\rho_{\alpha}^{\alpha+1}\left(\mathcal{I}\right)=0$.
For statement 2, we handle the cases when $\alpha=0$ and $\alpha\geq 1$ separately. If $\alpha=0$, then $\Lambda_{\alpha+1}=\Lambda_{1}$ is a directed graph, and as in Example 6.2, [6, Proposition 3.10] implies $\mathcal{O}\left(X_{0}\right)\cong\mathcal{O}\left(\Lambda_{1}\right)$. If $\alpha\geq 1$, then $J_{\alpha}=0$. Since $\left(\psi_{\alpha},\pi_{\alpha}\right)$ is a Toeplitz representation of $X_{\alpha}$, this implies $\left(\psi_{\alpha},\pi_{\alpha}\right)$ is Cuntz-Pimsner covariant. Covariance then induces a *-homomorphism $\psi_{\alpha}\times\pi_{\alpha}:\mathcal{O}\left(X_{\alpha}\right)\rightarrow\mathcal{O}\left(\Lambda_{\alpha+1}\right)$. We claim $\psi_{\alpha}\times\pi_{\alpha}$ is an inverse for $j_{\alpha}$. Proving this is simple since we may check the composition on generators. For $e\in\Lambda^{\omega^{\alpha}}$,
$\left(\psi_{\alpha}\times\pi_{\alpha}\circ j_{\alpha}\right)\left(T_{e}\right)=\left(\psi_{\alpha}\times\pi_{\alpha}\right)\left(\mu_{\alpha}\left(\chi_{e}\right)\right)=\psi_{\alpha}\left(\chi_{e}\right)=T_{e}$
Likewise for $e\in\Lambda_{\alpha}$,
$\left(\psi_{\alpha}\times\pi_{\alpha}\circ j_{\alpha}\right)\left(T_{e}\right)=\left(\psi_{\alpha}\times\pi_{\alpha}\right)\left(\eta_{\alpha}\left(T_{e}\right)\right)=\pi_{\alpha}\left(T_{e}\right)=T_{e}$
Thus $\psi_{\alpha}\times\pi_{\alpha}\circ j_{\alpha}=\text{id}$, and $j_{\alpha}$ is injective. Since $\mathcal{O}\left(X_{\alpha}\right)$ is generated by elements of the form $\mu_{\alpha}\left(\chi_{e}\right)$ and $\eta_{\alpha}\left(T_{e}\right)$, $j_{\alpha}$ is also surjective. Hence $j_{\alpha}$ is an isomorphism.
For statement 3, let $\mathcal{F}$ be the set of all connected components of $\Lambda_{\alpha}$. Note that $\rho_{\alpha}^{\alpha+1}\left(S_{v}\right)=T_{v}\not=0$ for $v\in\Lambda_{0}$, so $\rho_{\alpha}^{\alpha+1}$ is injective. Then by 3.25,
$\mathcal{A}_{\alpha}=\rho_{\alpha}^{\alpha+1}\left(\mathcal{O}\left(\Lambda_{\alpha}\right)\right)\cong\bigoplus_{F\in\mathcal{F}}\rho_{\alpha}^{\alpha+1}\left(\mathcal{O}\left(F\right)\right)$
Note that the operation above is the $c_{0}$ direct sum. For $a\in\mathcal{A}_{\alpha}$ with $a\geq 0$, $n\in\mathbb{N}$, and $\varepsilon>0$ we wish to show there exists $m\geq n$ and non-returning $\xi\in X_{\alpha}^{\otimes m}$ such that $\|\xi\|=1$ and
$\left\|\left\langle\xi,\varphi_{\alpha}\left(a\right)\xi\right\rangle_{\alpha}\right\|>\|a\|-\varepsilon$
We begin by letting $a\in\mathcal{A}_{\alpha}$, $a\geq 0$, $n\in\mathbb{N}$, and $\varepsilon>0$ be given. By the isomorphism above, choose $F\in\mathcal{F}$ such that $\left\|a\right\|=\left\|p_{F}\left(a\right)\right\|$, where $p_{F}:\mathcal{A}_{\alpha}\rightarrow\rho_{\alpha}^{\alpha+1}\left(\mathcal{O}\left(F\right)\right)$ projects onto the $F$-component of the direct sum. Since $\Lambda_{\alpha+1}$ satisfies condition (S), there exists $m\geq n$ and non-returning, $\alpha$-full $f\in\Lambda_{\alpha+1}$ such that $r\left(f\right)\in F$ and $\omega^{\alpha}\cdot m\leq d\left(f\right)<\omega^{\alpha}\cdot\left(m+1\right)$. Then by 7.9, $\chi_{f}\in X_{\alpha}^{\otimes m}$ is non-returning. Applying 7.10, we obtain the following closed Hilbert submodule $K$ of $X_{\alpha}^{\otimes m}$ consisting only of non-returning vectors.
$K=\overline{\text{$\mathrm{span}$}}\left\{\varphi_{\alpha}\left(a\right)\chi_{f}\cdot b:a,b\in\mathcal{A}_{\alpha}\right\}$
Since $K$ is closed under the left action $\varphi_{\alpha}$, $\varphi_{\alpha}$ restricts to a *-homomorphism $\tau:\rho_{\alpha}^{\alpha+1}\left(\mathcal{O}\left(F\right)\right)\rightarrow\mathcal{L}\left(K\right)$. Here we apply the $\alpha$-fullness of $f$ to show that for each $w\in\Lambda_{0}\cap F$, $\left(\tau\circ\rho_{\alpha}^{\alpha+1}\right)\left(S_{w}\right)=\tau\left(T_{w}\right)\not=0$. For $w\in\Lambda_{0}\cap F$, select $\gamma<\omega^{\alpha}$ and $e\in F$ such that $r\left(f^{\gamma}\right)=s\left(e\right)$ and $r\left(e\right)=w$. Then $\varphi_{\alpha}\left(T_{e}T_{f_{\gamma}}^{*}\right)\chi_{f}\in K$ is non-returning, and
$\tau\left(T_{w}\right)\left(\varphi_{\alpha}\left(T_{e}T_{f_{\gamma}}^{*}\right)\chi_{f}\right)=\tau\left(T_{w}\right)\left(\chi_{ef^{\gamma}}\right)=\chi_{ef^{\gamma}}$
Moreover, $\chi_{ef^{\gamma}}\not=0$ because
$\left(\mu_{\alpha}^{\otimes m}\circ j_{\alpha}^{-1}\right)\left(\chi_{ef^{\gamma}}\right)=T_{ef^{\gamma}}\not=0$
In particular, $\tau\left(T_{w}\right)\not=0$.
Since $F$ satisfies condition (S), the induction hypothesis implies $\tau$ is injective, and in particular isometric. Choose $b\in\mathcal{A}_{\alpha}$ such that $a=b^{*}b$. Then
$\left\|b\right\|=\left\|a\right\|^{1/2}=\left\|p_{F}\left(a\right)\right\|^{1/2}=\left\|p_{F}\left(b\right)\right\|=\left\|\left(\tau\circ p_{F}\right)\left(b\right)\right\|$
Using continuity of $t\mapsto t^{2}$, we choose $\delta>0$ such that $\left(\left\|b\right\|-\delta\right)^{2}>\left\|b\right\|^{2}-\varepsilon=\left\|a\right\|-\varepsilon$. Then we choose $\xi\in K$ such that $\left\|\xi\right\|=1$ and $\left\|\left(\tau\circ p_{F}\right)\left(b\right)\xi\right\|>\left\|b\right\|-\delta$. This implies
$\left\|\left\langle\xi,\varphi_{\alpha}\left(a\right)\xi\right\rangle_{\alpha}\right\|$ $=\left\|\left\langle\varphi_{\alpha}\left(b\right)\xi,\varphi_{\alpha}\left(b\right)\xi\right\rangle_{\alpha}\right\|$
$=\left\|\varphi_{\alpha}\left(b\right)\xi\right\|^{2}$
$=\left\|\left(\tau\circ p_{F}\right)\left(b\right)\xi\right\|^{2}$
$>\left(\left\|b\right\|-\delta\right)^{2}$
$>\left\|a\right\|-\varepsilon$
Thus $X_{\alpha}$ satisfies condition (S).
Now we prove that $\mathcal{O}\left(\Lambda_{\alpha+1}\right)\cong\mathcal{O}\left(X_{\alpha}\right)$ has Cuntz-Krieger uniqueness. Let $\pi:\mathcal{O}\left(\Lambda_{\alpha+1}\right)\rightarrow\mathcal{C}$ be a *-homomorphism into a $\mathrm{C}^{*}$-algebra $\mathcal{C}$ such that $\pi\left(T_{v}\right)\not=0$ for all $v\in\Lambda_{0}$. Then $\left(\pi\circ\rho_{\alpha}^{\alpha+1}\right)\left(S_{v}\right)\not=0$ for each $v\in\Lambda_{0}$. Since $\mathcal{O}\left(\Lambda_{\alpha}\right)$ has Cuntz-Krieger uniqueness by assumption, $\pi\circ\rho_{\alpha}^{\alpha+1}$ is injective, and in particular, $\left.\pi\right|_{\mathcal{A}_{\alpha}}$ is injective. Because $\psi_{\alpha}\times\pi_{\alpha}:\mathcal{O}\left(X_{\alpha}\right)\rightarrow\mathcal{O}\left(\Lambda_{\alpha+1}\right)$ is an isomorphism, $\left.\left(\pi\circ\psi_{\alpha}\times\pi_{\alpha}\right)\right|_{\eta_{\alpha}\left(\mathcal{A}_{\alpha}\right)}$ is injective. Applying [4, Theorem 3.9], we see $\pi$ is injective.
Finally, we must show that if $\beta$ is a limit ordinal and for all $\gamma<\beta$, $\mathcal{O}\left(\Lambda_{\gamma}\right)$ has Cuntz-Krieger uniqueness, then $\mathcal{O}\left(\Lambda_{\beta}\right)$ has Cuntz-Krieger uniqueness. Let $\left\{V_{e}:e\in\Lambda\right\}$ be the generators of $\mathcal{O}\left(\Lambda_{\beta}\right)$ and $\pi:\mathcal{O}\left(\Lambda_{\beta}\right)\rightarrow\mathcal{C}$ be a *-homomorphism with $\pi\left(V_{v}\right)\not=0$ for all $v\in\Lambda_{0}$. By 3.23, we have the following
$\mathcal{O}\left(\Lambda_{\beta}\right)=\overline{\bigcup_{\gamma<\beta}\rho_{\gamma}^{\beta}\left(\mathcal{O}\left(\Lambda_{\gamma}\right)\right)}$
For $\gamma<\beta$, $\pi\circ\rho_{\gamma}^{\beta}$ is injective, hence $\ker\pi\cap\rho_{\gamma}^{\beta}\left(\mathcal{O}\left(\Lambda_{\gamma}\right)\right)=0$. By [1, II.8.2.4], we get $\ker\pi=0$, and $\pi$ is injective. ∎
8. Condition (S)
Given an arbitrary ordinal graph $\Lambda$, it’s not clear how to verify condition (S) in 7.13. Our final goal is to derive a sufficient condition for condition (S) given an ordinal graph which satsifies condition (V) defined below. We prove that it suffices to check a class of directed graphs $\left\{\mathcal{F}_{\alpha}:\alpha\in\mathrm{Ord}\right\}$. Throughout this section, let $\Lambda$ be a fixed ordinal graph. Recall that $e_{\alpha}$ and $e^{\alpha}$ defined in 3.9 are the unique paths for which $d\left(e_{\alpha}\right)=\alpha$ and $e=e_{\alpha}e^{\alpha}$.
Definition 8.1.
$\Lambda$ satisfies condition (V) if for every $\alpha\in\mathrm{Ord}$ and $e\in\Lambda^{\omega^{\alpha}}$, $e$ is $\alpha$-full.
Remark 8.2.
Since every path in $\Lambda^{\omega^{0}}=\Lambda^{1}$ is 00-full, every directed graph satisfies condition (V).
Definition 8.3.
For $\alpha\in\mathrm{Ord}$ and $f,g\in\Lambda^{\omega^{\alpha}}$, let $f\cong_{\alpha}g$ if and only if there exist $\beta,\gamma<\omega^{\alpha}$ such that $f^{\beta}=g^{\gamma}$.
Clearly the relation $\cong_{\alpha}$ is symmetric. It is also reflexive since we may choose $\beta=\gamma=0$. Additionally, if $f,g,h\in\Lambda^{\omega^{\alpha}}$ with $f\cong_{\alpha}g$ and $g\cong_{\alpha}h$, choose $\beta,\gamma,\delta,\epsilon<\omega^{\alpha}$ such that $f^{\beta}=g^{\gamma}$ and $g^{\delta}=h^{\epsilon}$. Then without loss of generality, assume $\gamma\leq\delta$, in which case
$f^{\beta-\gamma+\delta}=\left(f^{\beta}\right)^{-\gamma+\delta}=\left(g^{\gamma}\right)^{-\gamma+\delta}=g^{\gamma-\gamma+\delta}=g^{\delta}=h^{\epsilon}$
Hence $f\cong_{\alpha}h$. Thus $\cong_{\alpha}$ is an equivalence relation.
Recall that for each ordinal graph $\Gamma$ we have an equivalence relation $\sim$ from 3.24 such that for $e,f\in\Gamma$, $e\sim f$ if and only if $e$ and $f$ belong to the same connected component of $\Gamma$. In particular, we have one such relation $\sim_{\alpha}$ for each $\alpha\in\mathrm{Ord}$ by setting $\Gamma=\Lambda_{\alpha}$.
Definition 8.4.
For $\alpha\in\mathrm{Ord}$, define a directed graph $\mathcal{F}_{\alpha}=\left(\mathcal{F}_{\alpha}^{0},\mathcal{F}_{\alpha}^{1},r_{\alpha},s_{\alpha}\right)$ with vertices $\mathcal{F}_{\alpha}^{0}=\Lambda_{\alpha}/\sim_{\alpha}$ and edges $\mathcal{F}_{\alpha}^{1}=\Lambda^{\omega^{\alpha}}/\cong_{\alpha}$. Define the range and source maps by
$r_{\alpha}\left(\left[e\right]_{\cong_{\alpha}}\right)$ $=\left[r\left(e\right)\right]_{\sim_{\alpha}}$
$s_{\alpha}\left(\left[e\right]_{\cong_{\alpha}}\right)$ $=\left[s\left(e\right)\right]_{\sim_{\alpha}}$
If $\beta,\gamma<\omega^{\alpha}$ with $e^{\beta}=f^{\gamma}$, then $s\left(e\right)=s\left(e^{\beta}\right)=s\left(f^{\gamma}\right)=s\left(f\right)$. Thus the source map $s_{\alpha}$ of $\mathcal{F}_{\alpha}$ is well-defined. It remains to see the range map is well-defined. Suppose $e^{\beta}=f^{\gamma}$. Then $d\left(e_{\beta}\right)=\beta<\omega^{\alpha}$, so $s\left(e_{\beta}\right)=r\left(e^{\beta}\right)\sim_{\alpha}r\left(e_{\beta}\right)=r\left(e\right)$. For the same reason, $r\left(f^{\gamma}\right)\sim_{\alpha}r\left(f\right)$. Since $r\left(e^{\beta}\right)=r\left(f^{\gamma}\right)$, we see $r\left(e\right)\sim_{\alpha}r\left(f\right)$. Thus the range map $r_{\alpha}$ is well-defined.
We analyze $\Lambda$ in terms of entries of cycles in the directed graphs $\mathcal{F}_{\alpha}$. We use the definitions of cycle and entry from [9], which we record below.
Definition 8.5 ([9, pg. 16]).
If $E=\left(E^{0},E^{1},r_{E},s_{E}\right)$ is a directed graph, a cycle is a path $\mu=\mu_{1}\mu_{2}\ldots\mu_{n}$ with $n\geq 1$, $\mu_{k}\in E^{1}$, $s_{E}\left(\mu_{n}\right)=r_{E}\left(\mu_{1}\right)$ and $s_{E}\left(\mu_{j}\right)\not=s_{E}\left(\mu_{k}\right)$ for $j\not=k$. An entry to a cycle $\mu$ in $E$ is an edge $e\in E^{1}$ such that there is $j$ with $r_{E}\left(\mu_{j}\right)=r_{E}\left(e\right)$ and $\mu_{j}\not=e$.
Definition 8.6 ([9, Lemma 3.7]).
If $E$ is a directed graph and $e=e_{1}e_{2}\ldots e_{n}$ is a path in $E$ with $e_{j}\in E^{1}$, then $e$ is non-returning if $e_{j}\not=e_{n}$ for $j<n$.
Note that this definition of non-returning is slightly different than the definition of a non-returning path in an ordinal graph from 7.2. In the following proof, we apply the above definition for the directed graphs $\mathcal{F}_{\alpha}$, since it is more convenient, and use 7.2 for ordinal graphs. Raeburn proves in [9, Lemma 3.7] that if every cycle of a directed graph has an entry, then for every $n\in\mathbb{N}$ and $v\in E^{0}$ there exists a non-returning path $e$ with $r\left(e\right)=v$ whose length is at least $n$. We use this in the proof of the following theorem.
Theorem 8.7.
Suppose $\Lambda$ satisfies condition (V) and for some $\alpha\in\mathrm{Ord}$, $\Lambda_{\alpha}$ satisfies condition (S). If every cycle in $\mathcal{F}_{\alpha}$ has an entry, then $\Lambda_{\alpha+1}$ satisfies condition (S).
Proof.
We must show for every $v\in\Lambda_{0}$ and $n\in\mathrm{Ord}$ with $n<\omega$, there exists $k$ with $\omega>k\geq n$ and $u\in\Lambda^{\omega^{\alpha}\cdot k}$ such that $r\left(u\right)=v$ and $u$ is non-returning. Since every cycle of $\mathcal{F}_{\alpha}$ has an entry, choose $k\geq n$ and a non-returning path $z=f_{1}f_{2}f_{3}\ldots f_{k}$ in $\mathcal{F}_{\alpha}$ with $r_{\alpha}\left(z\right)=\left[v\right]_{\sim_{\alpha}}$ and $f_{j}\in\mathcal{F}_{\alpha}^{1}$. For each $1\leq j\leq k$, choose $g_{j}\in\Lambda^{\omega^{\alpha}}$ such that $\left[g_{j}\right]_{\cong_{\alpha}}=f_{j}$. Then $\left[r\left(g_{1}\right)\right]_{\sim_{\alpha}}=r_{\alpha}\left(f_{j}\right)=r_{\alpha}\left(z\right)=\left[v\right]_{\sim_{\alpha}}$, hence $r\left(g_{1}\right)\sim_{\alpha}v$. Since $\Lambda$ satisfies condition (V), there exists $\beta_{1}<\omega^{\alpha}$ and $h_{1}\in\Lambda_{\alpha}$ such that $s\left(h_{1}\right)=r\left(g_{1}^{\beta_{1}}\right)$ and $r\left(h_{1}\right)=v$. Similarly, $r\left(g_{2}\right)\sim_{\alpha}s\left(g_{1}\right)$, so we may choose $\beta_{2}<\omega^{\alpha}$ and $h_{2}\in\Lambda_{\alpha}$ such that $s\left(h_{2}\right)=r\left(g_{2}^{\beta_{2}}\right)$ and $r\left(h_{2}\right)=s\left(g_{1}\right)$. Repeat this process for $2<j\leq k$ to construct $h_{j}$ and $\beta_{j}<\omega^{\alpha}$ such that $s\left(h_{j}\right)=r\left(g_{j}^{\beta_{j}}\right)$ and $r\left(h_{j}\right)=s\left(g_{j-1}\right)$. For each $j$ define $p_{j}=h_{j}g_{j}^{\beta_{j}}$.
Since $p_{j}^{d\left(h_{j}\right)}=\left(h_{j}g_{j}^{\beta_{j}}\right)^{d\left(h_{j}\right)}=g_{j}^{\beta_{j}}$, $\left[p_{j}\right]_{\cong_{\alpha}}=f_{j}$. For $j<k$, $s\left(p_{j}\right)=s\left(g_{j}\right)=r\left(h_{j+1}\right)$. Hence we may define $u=p_{1}p_{2}\ldots p_{k}$. Then $r\left(u\right)=r\left(p_{1}\right)=v$, and because $d\left(h_{j}\right)<\omega^{\alpha}$ and $\beta_{j}<\omega^{\alpha}$,
$d\left(p_{j}\right)=d\left(h_{j}\right)+d\left(g_{j}^{\beta_{j}}\right)=d\left(h_{j}\right)-\beta_{j}+d\left(g_{j}\right)=d\left(h_{j}\right)-\beta_{j}+\omega^{\alpha}=\omega^{\alpha}$
Thus $d\left(u\right)=\omega^{\alpha}\cdot k$. We claim $u$ is non-returning in the sense of 7.2. Towards a contradiction, suppose $u$ is not non-returning. Then there exists $e\in\Lambda$ with $\omega^{\alpha}\leq d\left(e\right)<\omega^{\alpha}\cdot k$ and $\beta<\omega^{\alpha}$ such that $eu^{\beta}\in u\Lambda$. Let $q\in\Lambda$ with $eu^{\beta}=uq$. Choose $m<k$ such that $\omega^{\alpha}\cdot m\leq d\left(e\right)<\omega^{\alpha}\cdot\left(m+1\right)$ and $\gamma<\omega^{\alpha}$ with $d\left(e\right)=\omega^{\alpha}\cdot m+\gamma$. Choose $x_{1},x_{2},\ldots,x_{m}\in\Lambda^{\omega^{\alpha}}$ and $t\in\Lambda_{\alpha}$ such that $e=x_{1}x_{2}\ldots x_{m}t$. Then $u^{\beta}=\left(p_{1}\right)^{\beta}p_{2}\ldots p_{k}$, hence
$x_{1}x_{2}\ldots x_{m}t\left(p_{1}\right)^{\beta}p_{2}\ldots p_{k}=p_{1}p_{2}\ldots p_{k}q$
Now we split into two cases. In the first case, we have $k=m+1$. Then by unique factorization,
$t\left(p_{1}\right)^{\beta}p_{2}\ldots p_{k}=p_{k}q$
Then $d\left(t\left(p_{1}\right)^{\beta}\right)=d\left(t\right)-\beta+d\left(p_{1}\right)=d\left(t\right)-\beta+\omega^{\alpha}=\omega^{\alpha}$. Thus
$\left(p_{k}q\right)_{\omega^{\alpha}}=p_{k}=t\left(p_{1}\right)^{\beta}$
Therefore $\left(p_{k}\right)^{d\left(t\right)}=\left(p_{1}\right)^{\beta}$, and $\left[p_{k}\right]_{\cong_{\alpha}}=f_{k}=\left[p_{1}\right]_{\cong_{\alpha}}=f_{1}$. Since $z$ is a non-returning path in $\mathcal{F}_{\alpha}$, this is a contradiction. For the other case, suppose $k>m+1$. Similarly, unique factorization implies
$p_{k-m}p_{k-m+1}\ldots p_{k}=p_{k}q$
Thus $p_{k-m}=p_{k}$ and $f_{k-m}=f_{k}$, again contradicting the fact that $z$ is non-returning. ∎
Remark 8.8.
If $\beta$ is a limit ordinal and $\Lambda_{\alpha}$ satisfies condition (S) for every $\alpha<\beta$, then $\Lambda_{\beta}$ satisfies condition (S). Thus 8.7 tells us that to verify $\Lambda$ satisfies condition (S), it suffices to check $\Lambda$ satisfies condition (V) and for every $\alpha\in\mathrm{Ord}$, every cycle of $\mathcal{F}_{\alpha}$ has an entry.
Example 8.9.
Consider the ordinal graph in Fig. 8.1 generated by one vertex, two paths of length 1, and two paths of length $\omega$. $\mathcal{O}\left(\Lambda\right)$ is the $\mathrm{C}^{*}$-algebra generated by two isometries $T_{e},T_{f}$ with $T_{e}T_{e}^{*}+T_{f}T_{f}^{*}=T_{v}=1$ and two isometries $T_{g},T_{h}$ with $T_{e}T_{g}=T_{g}$ and $T_{f}T_{h}=T_{h}$. Then $\mathcal{F}_{0}$ and $\mathcal{F}_{1}$ each are directed graphs with one vertex and two edges. $\Lambda$ satisfies condition (V) since it has only one vertex, and every cycle in $\mathcal{F}_{0}$ and $\mathcal{F}_{1}$ has an entry. Therefore $\Lambda$ satisfies condition (S). Because $\Lambda$ has no 1-sources and no 1-regular vertices, 7.15 implies $\mathcal{O}\left(\Lambda\right)$ has Cuntz-Krieger uniqueness. If $\pi:\mathcal{O}\left(\Lambda\right)\rightarrow\mathcal{C}$ is a *-homomorphism, then either $\pi$ is injective or $T_{v}\in\ker\pi$. But $T_{v}$ is a unit for $\mathcal{O}\left(\Lambda\right)$, so in this case $\ker\pi=\mathcal{O}\left(\Lambda\right)$. Thus $\mathcal{O}\left(\Lambda\right)$ is simple.
$v$$e$$f$ $g:eee\dots$ $g=eg$ $h:fff\dots$ $h=fh$
Figure 8.1. The ordinal graph $\Lambda$ in Example 8.9
We would like to end with a more complicated example. The previous examples of ordinal graphs all have path lengths which are bounded by $\omega^{2}$. Now we introduce an example of an ordinal graph whose path lengths are bounded by $\omega^{\omega}$.
Example 8.10.
Consider the directed graph in Fig. 8.2. We wish to construct an ordinal graph $\Lambda$ in which for every $\alpha\in\mathbb{N}$, the connected components of $\mathcal{F}_{\alpha}$ are isomorphic to this graph. We consider an inductive construction, starting with $\Gamma_{1}=\mathcal{F}_{0}$ as this directed graph. Then to get $\Gamma_{2}$, we duplicate the previous directed graph to get both vertices of $\mathcal{F}_{1}$. We then add a path of length $\omega$ for each edge in $\mathcal{F}_{1}$. Duplicating this ordinal graph again gives us both vertices of $\mathcal{F}_{2}$, and adding paths of length $\omega^{2}$ gives us each edge in $\mathcal{F}_{2}$. Regarding at each step $\Gamma_{k}$ as a subgraph of the ordinal graph $\Gamma_{k+1}$ and taking the union $\cup_{k\in\mathbb{N}}\Gamma_{k}$, we obtain an ordinal graph $\Lambda$ whose path lengths are bounded by $\omega^{\omega}$. Of course, this construction is neither complete nor formal, as we have not specified how these paths factor. Below we formally define one choice for $\Lambda$ using generators and relations.
Define $X=\left\{x\in\left\{0,1\right\}^{\mathbb{N}}:x^{-1}\left(1\right)\text{ is finite}\right\}$. For $x\in X$, define $0x\in X$ and $1x\in X$ by
$\left(0x\right)\left(n\right)$ =$=\begin{cases}x\left(n-1\right)&n\geq 1\\ 0&n=0\end{cases}$
$\left(1x\right)\left(n\right)$ =$=\begin{cases}x\left(n-1\right)&n\geq 1\\ 1&n=0\end{cases}$
Let $\Lambda$ be the category generated by $\left\{v_{x},e_{x}^{\alpha},f_{x}^{\alpha},g_{x}^{\alpha},h_{x}^{\alpha}:x\in X,\alpha\in\mathbb{N}\right\}$ with the following relations:
  1. (1)
    $v_{0x}=s\left(e_{x}^{0}\right)=r\left(e_{x}^{0}\right)=s\left(g_{x}^{0}\right)=r\left(h_{x}^{0}\right)$
  2. (2)
    $v_{1x}=s\left(f_{x}^{0}\right)=r\left(f_{x}^{0}\right)=r\left(g_{x}^{0}\right)=s\left(h_{x}^{0}\right)$
  3. (3)
    $s\left(e_{x}^{\alpha+1}\right)=s\left(e_{0x}^{\alpha}\right)$
  4. (4)
    $s\left(f_{x}^{\alpha+1}\right)=s\left(f_{1x}^{\alpha}\right)$
  5. (5)
    $s\left(g_{x}^{\alpha+1}\right)=s\left(g_{0x}^{\alpha}\right)$
  6. (6)
    $s\left(h_{x}^{\alpha+1}\right)=s\left(h_{1x}^{\alpha}\right)$
  7. (7)
    $e_{0x}^{\alpha}e_{x}^{\alpha+1}=e_{x}^{\alpha+1}$
  8. (8)
    $f_{1x}^{\alpha}f_{x}^{\alpha+1}=f_{x}^{\alpha+1}$
  9. (9)
    $g_{1x}^{\alpha}h_{1x}^{\alpha}g_{x}^{\alpha+1}=g_{x}^{\alpha+1}$
  10. (10)
    $h_{0x}^{\alpha}g_{0x}^{\alpha}h_{x}^{\alpha+1}=h_{x}^{\alpha+1}$
There is a length functor $d:\Lambda\rightarrow\mathrm{Ord}$ defined by
$d\left(v_{x}\right)$ $=0$
$d\left(e_{x}^{\alpha}\right)$ $=d\left(f_{x}^{\alpha}\right)=d\left(g_{x}^{\alpha}\right)=d\left(h_{x}^{\alpha}\right)=\omega^{\alpha}$
This makes $\Lambda$ an ordinal graph. Indeed, for each $\alpha\in\mathbb{N}$ the (infinitely many) connected components of $\mathcal{F}_{\alpha}$ are isomorphic to the graph in Fig. 8.2. $\Lambda$ satisfies condition (V), so by 8.7, $\Lambda$ satisfies condition (S).
$e$$f$$g$$h$
Figure 8.2. A connected component of $\mathcal{F}_{\alpha}$ for $\Lambda$ defined in Example 8.10